A Complete Bibliography of Publications in Obituary Notices of Fellows of the Royal

Society

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.03

Title word cross-reference

17 [Mil45a]. 1928 [B.33c, Gib38]. 1930 [B32b, H32a, W.32a, W.32b]. 1931 [B.32a, F.32b, F32a, J33a, N.32]. 1932 [A.32, B.32c, B.32d, B.33b, C.33, D.32b, D.32a, D’E32, GP32, H.32b, H.33b, R.32, R.33b, R.33a, Syn33]. 1933 [Ano34b, Ano34c, Ano34d, Ano34e, B.33a, D.34a, E.33, G.34, Gar34, H.33c, H.33a, HS34, J.33b, K.33, MB33, M.34c, M.34a, N.33, Wil34]. 1934 [Ano34a, Con35, D.34b, Dou35b, Dou35a, FT35, Hil35a, L.34a, L.34b, LG35, M.34b, Maw35, Mil35, Nut35, Pop35, R.34, S.34b, S.34a, She35, Sim35, T.34, W.34]. 1935 [Arm35, Bri36, Car36, Cat35, Dal35, Ell35, Eve35, Gla35, Hal35, Har35, Hil35b, Ken35, Pou35, Rap35, RS36, Tho35, Whi35, Woo36]. 1936 [Anr36, D’E36, Dav36, Dou36, Fea36, Hic36, Ing36, Lai36, Par36, Rob36, Spe36, Whi36, YF36]. 1937 [And38, Atk38, Bai39, Bak38, Con39, EC38, GSK39, Gre39b, Imm38, Jef39, Jor38, Kee41, Kei38, Riv39, Rob39, Sah40, Smi38, Smi39, Sti39, Sum38, Tan39, Tho38, Tur38, Twy38, Wat40a, Wil38]. 1938 [Ano39, Dal39, Dar40, Des38, Dys39, Fea38, Fin39, FT38, Gib40, Gre39a, GS40, Hil39, Hol39, Hop38, Jon38, Lar38, Mar38, Mar39, Moo38, Mur38, New39, Pop38, Pra39, Sod39, Tho39, Wen38, Whi39, WW38]. 1939 [Can41, Cla40, D’E40, Dal40, Edd40, Ell41, Gib41, WGS40, Had40, Jac47, Lea40, Mor40, Ran40, Rap40, ST40, Ste40, Tan41, Wat40b, Whi40]. 1940 [Bra41, Cal41, Dal41, Des41, Din41, Edw41, Fle41a, Gar41a, GF41, Had41, Hol41, HM42, Imm41, Ing41, Irv41, Jon41, Mic41, Mil41, Mit41, Mur41, Mye41, Rus41, Sou41, Str41, Tho41b, Til41, Tur41, Whi41]. 1941 [Bad43, Bes42, Bro42, Bur44b, Car42, Don42, Ege42, Fle41b, Gar41b, Gra42, Gri41, Guy41, Had42, Har41, Hil41, Hod42, Jon42, Jon43, LD41, LS42, Myr41, Rap41, Rid42, Rob47, Rot41, Smi41, Tho41a, VB41]. 1942 [AL43, Bal43, Dal42, Dow42, Edd42, Fle43, Har43, Har54, Har42, HS43, Kei43, KH43, Ker43, Rob53, Rus42, Sim44, Smi45, Tow43, Whi42]. 1943 [Bai43, Bak44, Bal44, BN43, Cam43, Car44, Dam44, Des44, Ege43, FCH44, Gra44, Gre43, LC44, Mas43, Rou47, Rus44, Str44b, Wen44, Wey44]. 1944 [All47, Bed47, Bla45, Bro45, Bur44a, Cle44, Coo45, Dal44, Don48b, Gra45, Gre44, Kei44, Kin44, Mar45, Mil44b, Mil45a, Mil45b, Plu45, Rid44, Rob45, Str44a]. 1945 [AS45, Bla47, Cal47, Dal47, Des45, DG45, Ecc45, Ell45, Fer45, Fis47, GH48, Gol45, Hev48, Hil49, Hin47, JF48, Lan47, Rus48, Tho47, dB47]. 1946 [Bar48, Bee47, Chr47a, Chr47b, Erd47, FC47, Gar48, Gre48, Hil47, Jes53, Jol51, Kei48, Mil47, Pic48, Ric48, Rod48, Rou49b, Toi48]. 1947 [Bor48, Bos48, Bri48, Bro48, Col48, Dal48, Ege49, Ell48, Fer48, Haw48, Hin48, Irv48, MC48, Mun53, PM49, Rea48, Rob48, Rou49a, Til48, Tit49, Wat50, Whi48]. 1948 [CG48, Dob49, Don48a, Ede50, Fos49, Gre51, Har50, Hau49, Hoa49, Lan49, De 52, MW48, Rob49, Ryd49, Sch50, Str54, Tay49, Tho49]. 1949 [Ano50a, Cha50, Eng50, Far49, Ham50, Hil50, Hog50, Ken51, Ken50, Kin50, Lew50, Mur50, Par50, Smi50, Whi49, Wig49]. 1950 [And52a, Ben51, Cal51, Cam51, DKY52, Edg51, Fil51, Fin51, Fos52, Hir51, Mas52, McC51, Rod51, Rus51, Sim50, Sut51, Whi51]. 1951 [Bor52, Fre54, Gol51, Har53, Lil51, Lin52, Pet54, Rea52, Rob52, Sal52, Tie52, Wat52, Whi52, Yat52]. 1952 [And52b, And53, Ano52, Bak53, Bla53, Bul53, Dal53, Dou53, Gan53, Gib52, GH53, Har52, Hop53, Lid52, Mel53, Moo53, Pip52, Rea53, Ros53, Str53, Tay54, Tiz54, Wil53, You54]. 1953 [Atk54, Bai54, Baw54, Ben53, Gar54, Gol54, Kei54, Rid53, Rus54, Tho53a, Tho53b, Whi54]. 1954 [Sal54, Wis54].

940 [Gar41c].

Abbott [Des41]. Abel [Dal39]. Abraham [N.32]. Adam [Par50]. Adams [Fle43]. Agar [Tie52]. Albert [Bai39, Hil35b, M.34a, Mur41, Wat40a, H32a, N.32, Tho41a]. Alcock [MB33]. Alexander [D.32a, Fin51, G.34, HS43, Kei38, Cam43, Hau49, Rea52, Rob48, Rus48].

Alexandre [Str54]. Alfred [D’E32, Gla35, De 52, Ano39, Din41, Fle41a, Gar41a, Lar38, MB33, Pra39, R.33b, Ran40, Rus42, ST40, Sti39, Tur41, VB41, Whi36, Whi48]. Allen [Gre43, KH43, Whi40]. Allmand [Fre54]. Almroth [Col48]. Ambrose [Ecc45]. Andrew [Whi42]. Andrewes [GP32]. Antoine [De 52]. Archibald [F.32b, Hop38, L.34a]. Arden [GH53]. Arden-Close [GH53]. Arkwright [Mar45]. Arman [Str53]. Armand [Lin52]. Armstrong [GH48, Kee41]. Arnold [Bor52]. Arnulph [B.33a]. Arthur [Fin39, Jon38, L.34b, Lar38, MC48, Mar45, McC51, Mit41, Sim35, Ano34a, Bad43, Bro42, Bro45, Coo45, Edg51, Fos49, Fre54, Gar41b, HM42, JF48, Ker43, Mar39, Myr41, Plu45, Rap40, Rob39, Rob47, Smi39, Tan39, Toi48, Wat40b, Whi51]. Ashworth [Hic36]. Aston [Hev48]. August [Bri36, Hil50]. Augustus [Pou35, Mil41, Wig49]. Ausin [Bro48].

Babkin [DKY52]. Bagnall [Car44]. Bailey [Tur38]. Baker [Tho35]. Balcarres [Bra41]. Baldwin [Irv48]. Balfour [Had40]. Baly [Don48a]. Banting [Bes42]. Baptiste [Tow43]. Barcroft [Rou49a]. Barger [Dal40]. Barkla [All47]. Barlow [Ell45, Pop35]. Barnard [Mur50]. Barnes [Fos52, Whi54]. Baron [Smi41]. Baronet [B.32d]. Barr [F.32b]. Barrois [Jac47]. Bart [Ell35]. Barton [Pra39]. Basil [Gib40]. Bate [HS34]. Bateman [Erd47]. Bather [L.34b]. Batten [Ell41]. Bedford [Mit41]. Beecher [Mor40]. Beit [B32b]. Bell [Rus41]. Bengough [Des45]. Benham [Ben51]. Bennett [Bak44]. Berend [Rap35]. Berkeley [Har42]. Bernard [Fea38]. Bertram [M.34b]. Bewdley [Irv48]. Biffen [Eng50]. Bijl [Sch50]. Birbal [Ham50]. Bjerknes [Gol51]. Black [S.34b]. Blackie [Gib38]. Blackman [Bri48]. Blaxland [Ben51]. Bolton [Ell48]. Bone [Fin39]. Boris [DKY52]. Bose [Sah40]. Boulenger [Wat40a]. Bourne [Gar41a, H.33c]. Bousfield [Dam44]. Bower [Lan49]. Boycott [Mar39]. Boys [Str44a]. Brachet [H32a]. Bradfor [Ell35]. Bradford [Dal47]. Bragg [AL43]. Brereton [Tho35]. Bretland [Bla45]. Brögger [Til41]. Brooks [Moo53]. Broom [Wat52]. Brown [Dar40, Gre51]. Browne [Hol39]. Bruce [B.32a, Smi50]. Bryan [B.33c, Bro48]. Bryce [Kei48]. Buller [Bro45]. Bulloch [LD41]. Burne [Kei54]. Burrard [LC44]. Butler [Mas43]. Byron [L.34a].

Cabourn [Ros53]. Cadman [Smi41]. Cajal [She35]. Calman [Gan53]. Calmette [M.34a]. Campbell [Dys39]. Cannon [Dal47]. Capel [And38]. Cardew [Whi36]. Carl [Tho38]. Carles [Jac47]. Carlton [Gre44]. Carlyle [Car36]. Carpenter [Edw41]. Cartan [Whi52]. Carter [Hop53]. Cash [Mar38]. Cathcart [Wis54]. Cavendish [Whi39]. Cawthorne [Wil34]. Cecil [Tay54, You54]. Chalmers [Hin47]. Chamberlain [Had41]. Chapman [R.33b]. Charles [Bed47, Don48a, H.33c, M.34a, Syn33, Tho41a, All47, Arm35, Bak53, Bar48, BN43, Dal53, Ell48, Fer45, GH53, Hoa49, Lid52, Mye41, Sim50, Sou41, Str44a, Tho47, Wil53, dB47]. Chaston [R.33b].

Chatelier [Des38]. Chattaway [Cle44]. Chattock [Ano34a]. Cheyne [B.32d]. Christiaan [And52a]. Christian [Bri36, Whi39]. Christopher [Til41]. Chunder [Sah40]. Church [Tan39]. Civita [Hod42]. Claridge [R.32]. Clark [VB41]. Clarkson [Gri41]. Claud [Ede50]. Clerk [R.33a]. Clifford [Ano50a, Ryd49]. Clinton [Gra44]. Clive [Bal44, Wat50]. Close [GH53]. Cockayne [Hil35a]. Cohen [Don48b, Rap35]. Coker [Jes53]. Coleman [Wat40b]. Coleridge [Gra44]. Collie [Bal43]. Colquhoun [Rea53]. Comrie [Mas52]. Conway [Whi51]. Conwy [Par36]. Cook [Gib52, Rob52]. Cooper [Wat50]. Copeman [MC48]. Copland [Ryd49]. Corrigenda [Ano44]. Corrigendum [Ano49]. Cort [Edw41, Fle41a]. Cossar [M.34c]. Cowell [Whi49]. Crawford [Bra41]. Cresswell [Gra42]. Crichton [Hol39]. Crompton [Rus41]. Cross [Arm35]. Crozier [Gre48]. Cruikshank [G.34]. Cunningham [Eve35]. Cushing [Can41]. Cuthbertson [Bal44]. Cyril [Don48a].

D’Abernon [Dal42]. Dakin [Har52]. Dalby [Ing36]. Dampier [Tay54]. Daniel [Cle44, Rus42, Wig49]. D’Arcy [Dob49]. David [B.32a, Maw35, A.32, Bra41, Bur44a, Wey44]. Davidson [S.34b]. Dawson [Fle43, WGS40]. Deslandres [Str54]. Devonshire [Whi39]. D’Eyncourt [Lil51]. Diller [W.32b]. Dillon [M.34b]. Dixey [Pou35]. Dixon [Atk54, Dav36, Whi36]. Dobell [Ano50a]. Douglas [Lai36]. Dr. [Fea38]. Dreyer [Dou35b]. Druce [R.32]. Drummond [You54]. Drysdale [Har52]. Dudley [Dal35]. Dugald [R.33a]. Duke [Mit41, Whi39]. Dukinfield [S.34a]. Dunstan [Des45, Ano50b]. Dye [A.32]. Dyson [Edd40].

E. [Jon41]. Eade [Tie52]. Earl [Bra41, Har42, Irv48]. Eddington [Plu45]. Edgar [Dal42, KH43, Mur38]. Edgeworth [Maw35]. Edmund [Kin50, Mor40]. Edward [Ano34c, Bak53, Col48, Hil35b, Hop38, Kee41, Pip52, Car44, CG48, Don48a, GH48, Gre43, Hil39, Kei43, McC51, Wis54]. Edwards [Imm41]. Edwin [Ano39, Mar39, Mur50, Bee47, Mas43, Tur38, W.32a]. Elford [And52b]. Elie [Whi52]. Ellcock [FCH44]. Ellery [New39]. Elliot [Wil38]. Elliott [Tur38]. Eltringham [Car42]. Emile [Had42, Ano34e]. Ernest [Ing36, Lan47, Rob36, Sod39, Tho53a, Ano34b, Ano52, Cal41, D.32b, Dar40, Jes53, Whi54]. Ernst [Bor48, Don48b]. Erskine [Dou53]. Eugene [Sal54]. Eustace [Smi38, Lil51]. Evan [Bla47]. Evans [Mel53, Myr41]. Eve [Fos49]. Evelyn [Rus41]. Ewart [And53, M.34c, Sti39]. Ewing [Gla35].

Fabry [dB47]. Fairfield [Woo36]. Falkiner [GSK39]. Faraday [Gar48]. Farmer [Ano52, Bla45]. Farr [Gra44]. Fawcett [Kei43]. Felix [Sal54]. Ferdinand [Bri36]. Fernandez [D’E32]. Field [Edg51, Ing41]. Fields [Syn33]. Filon [Jef39]. Finlay [W.34]. Fisher [Mur41]. Fleming [Ecc45]. Fletcher [E.33]. Flett [Rea48]. Flexner [Rou49b]. Flinders [Smi45]. Forbes [FT38]. Forster [AS45, Wat50]. Forster-Cooper [Wat50]. Forsyth [Whi42]. Fortescue [Ede50]. Fowler [Din41, Mil45b, Mil44a, Mil45a]. Fox [Rob45]. Francis [Bak38, D.32a, Hev48, L.34b, Par50, Sal52]. Francois [De 52]. Frank [Edd40, Mil44b, Ege43, Fle43, Kin44]. Frankland [Gar48, GH48]. Frazer [Fle41b]. Frederic [Cal51, Cha50]. Frederick [Arm35, GP32, GH53, Bes42, Bla53, Bri48, Cle44, Dal48, Fil51, Hop53, Imm41, Moo53, Pou35, Ric48, Rod48, Tho53a]. Fredrick [Lan49]. Freud [Tan41]. Freundlich [Don42]. Frewen [Sou41]. Friedrich [Bri36]. Friman [Gol51]. Fritsch [Sal54]. Frost [Bri48]. Fry [Fea36, Gol54]. Fryer [Ede50].

G [Con39]. Gabriel [Mye41]. Gardiner [FC47, Hil41]. Garrod [Hop38]. Garwood [Kin50]. Geer [Bai43]. Geoffrey [Bak44]. George [Bad43, D’E36, Dal44, Fle41b, Gre51, Jes53, Rob39, B.33c, Dal40, Dow42, FT38, GSK39, Gri41, Lew50, Lin52, New39, R.32, Sim44, Spe36, Wat40a, Yat52]. Georges [Dou35b]. Gerald [Dow42, Sim44]. Gerard [Bai43]. Gibbs [Gar41a]. Gibson [Jon43, Sim50]. Gilbert [H.33c, Hil47, Irv41, Rob52, Rus51]. Gilchrist [Car36]. Glauert [FT35]. Glazebrook [RS36]. Glover [All47]. Godfrey [Tit49]. Goebel [B.33b]. Goldschmidt [Til48]. Goodey [Baw54]. Goodrich [Bee47]. Goom [F32a]. Gordon [Gar54]. Gowland [Dal48]. Grafton [Wil38]. Graham [C.33]. Grant [Bed47, Bes42]. Green [Bad43]. Greenwood [Hog50]. Gregory [B.32c, Cam51, Str53]. Griffiths [D.32b]. Gulland [Haw48]. Guy [Des45, FCH44]. Gwynne [Mel53]. Gye [And53].

H. [Guy41]. Hadcock [D’E36]. Haddon [Fle41a]. Hadfield [Des41]. Haldane [Dou36]. Hale [New39]. Hall [Rus42]. Halliley [Dal53]. Harden [HM42]. Hardy [HS34, Tit49]. Harker [ST40]. Harmer [Cal51]. Harold [Ano52, Edw41, Tit49, Dal35]. Harry [Eng50, Tan39, Car42, Erd47, WGS40]. Hartley [B.33c, Hic36, Mur38]. Harvey [Can41]. Hastie [Kei48]. Hatfield [Des44]. Haworth [Hir51]. Head [Hol41]. Heath [Tho41b]. Hector [Whi35]. Heim [Bai39]. Hele [Guy41]. Hele-Shaw [Guy41]. Henderson [Lew50, Sim44]. Hendrik [Sch50]. Henri [Bur44b, Str54]. Henry [And38, AL43, Bro45, Cla40, Fer48, Gar54, GSK39, Har43, Har35, Mic41, S.34a, Tay49, Atk54, B.33a, Bul53, Dal44, Des38, Edw41, Gre48, Had40, Har52, Har53, HS43, Hol41, Kee41, Ros53, Wen38, Woo36]. Herbert [Des44, Gar34, Moo38, Rus54, Smi38, Whi49, Wil53, Cla40, Don42, Har35, J33a, MW48, Mur41, Tho35]. Herbrand [Mit41]. Herman [FT35]. Heron [Gre43]. Heron-Allen [Gre43]. Hicks [Mil35]. Hickson [Gar41c]. Higgins [Kei54]. Hilbert [Wey44]. Hill [Bro42, Dou53]. Hinks [JF48]. Hjort [Har50]. Hobson [Ano34b]. Holden [And38]. Holland [Fer48]. Holloway [Lan47]. Hopkins [Dal48]. Hopwood [Mil47]. Horace [LG35]. Horatio [Atk54]. Hough [Mil41]. Houston [G.34]. Howard [Ano39, D.32b, Mil44a, Mil45a, Mil45b, R.34, Fos52]. Howell [Pic48]. Howson [Ken50]. Hudson [Rus44]. Hugh [Lew50, Par50, Mil44b]. Hugo [Hal35]. Hunt [Fis47]. Hutchinson [Smi39].

Imms [Wig49]. Inglis [Bak53]. Innes [Hin48]. Irvine [Rea53]. Ivan [Anr36]. Ivanovitch [Har54].

Jack [You54]. Jackson [Gib41, Moo38]. Jacob [Bai43, Dal39, Rob45]. Jagadis [Sah40]. James [Bla47, Cat35, Chr47b, D.34a, Hol39, Ken35, Sti39, Ege42, Fin51, Fle41b, Gla35, Gre39b, H.33b, Hic36, Hil49, Kei44, M.34c, Mil47, Rea53, Rid53, Rus54, Sod39]. Jan [And52a]. January [Mil45a]. Jean [Tow43]. Jeans [Mil47]. Jenkin [Sou41]. Jocelyn [Ing41]. Johan [Har50]. Johannes [Bor52, Sch50]. John [B32b, D’E40, Ell35, Eve35, Fre54, Imm38, Mas43, Mas52, Myr41, Ano34c, Ano34d, B.32c, Bal43, Bed47, Bla45, Cat35, Chr47a, Dal39, Dou36, Ecc45, Ede50, FC47, Haw48, J.33b, Jon42, K.33, Kei38, Lan47, Lea40, Mar38, Mic41, Rap41, Rea48, Rid44, Rob45, Smi41, Syn33]. Johnson [KH43, Tho49, WW38]. Johnston [Kin50]. Joly [Ano34d]. Joseph [And52b, GF41, Whi52, Edd42, Gre39a, Mar45, Mur50, Rob36, Rou49a, Str41, Tay49]. Jost [PM49]. Julius [Don48b, Rap35]. July [Mil45a].

K.G [Irv48]. Karl [Bor48, B.33b, Rou47, YF36]. Kaye [Gri41]. Keeble [Bla53]. Keith [Rid44]. Kellaway [Dal53]. Kemp [Cal47]. Kendall [Fea36]. Kennedy [Gib38]. Kettle [Mur38]. Kipping [Cha50]. Kitchin [W.34]. Kon [Lin52]. Koren [Gol51]. Krogh [Hil50].

Laby [Pic48]. Lacroix [De 52]. Laidlaw [Dal41]. Lamb [LG35]. Lanchester [Ric48]. Landsteiner [Rou47]. Langevin [Jol51]. Lapworth [Rob47]. Larmor [Edd42]. Laurens [Mur41]. Lebesgue [Bur44b]. Ledingham [Bed47]. Lee [Kin44]. Lees [Wil53]. Léon [M.34a]. Leonard [Cam51, D.34a, Dou53, Hil35a]. Leslie [Mas52]. Levi [Hod42]. Levi-Civita [Hod42]. Lewis [DG45, Ell41, Hil47, Gol54]. Lindsay [Bra41]. Little [Tho41b]. Lloyd [Par36]. Lodge [GF41]. Lofft [And38]. Logie [Hau49]. Lord [Ege49, EC38, Jor38]. Lorimer [W.34]. Louis [Des38, Jef39]. Love [Mil41]. Lowry [Pop38]. Ludwig [Bor48, PM49]. Lugeon [Bai54]. Lusk [C.33]. Lyle [Gra45]. Lyons [Dal44].

Macallum [L.34a]. Macaulay [Bak38]. Macbride [Cal41]. MacDonald [Gre39b, Rap41, Whi35]. Maclagen [Tay49]. Macleod [Cat35]. MacWilliam [Kei38]. Major [Hog50]. Mallock [B.33a]. Marcus [Dou35a]. Marjory [Rob49]. Marr [Ano34c]. Marsh [Sod39]. Marshall [Par50]. Martin [Pop38, AS45]. Masson [Haw48, Riv39]. Masterman [Gar41b]. Mather [Sum38]. Matthew [Smi45, W.32b]. Maurice [Rod48, Bai54]. Max [Bor48, Tho38]. Maxwell [Smi38]. McBain [Rid53]. McDougall [GS40]. McFadden [Con35]. McKenzie [Rea52]. McLennan [Eve35]. McSwiney [Bro48]. Medforth [WGS40]. Mellanby [Lea40]. Mellor [Gre39a]. Mercer [H.33b]. Meredith [Mel53]. Mervyn [Gar54]. Meyerhof [Pet54]. Meyrick [Hil39]. Michelson [N.32]. Middlemiss [Fer45]. Middleton [Rus44]. Miers [HS43]. Miles [Rot41]. Millar [J.33b]. Milne [McC51]. Mitchell [Hin47]. Mitchinson [Mil35]. Moir [Kei44]. Monckton [MC48]. Mond [Tho39]. Morgan [Fis47, Irv41, Par36]. Moritz [Til48]. Morley [E.33, Hoa49]. Mostyn [Edg51]. Mott [Gib40]. Mowbray [Har42]. Muir [T.34]. Munro [Whi35]. Murray [Fin51]. Myers [Bar48].

N [Jon41]. Napier [Gol45]. Napoleon [Jef39]. Nernst [LS42]. Nevil [Tiz54]. Nevill [Jon41]. Neville [Had41]. Newall [Mil44b]. Newstead [Mun53]. Newton [Hil47, W.32a]. Nicholaevitch [Tho53b]. Nicolai [Har54]. Norman [Bal43, Hir51]. North [Whi48]. Notices [Mil45a]. Nuttall [GSK39].

Obituary [Mil45a]. Oldham [Dav36]. Oliver [Sal52, GF41]. Onslow [AS45]. Oram [D’E40]. Orme [Riv39]. Orpen [Lan49]. Orr [Con35]. Orton [Rus54]. Osborn [Woo36]. Otto [B32b, H.33a, Pet54].

Palmer [Rod51]. Parks [Jon38]. Parsons [Cam51]. Paterson [Ryd49]. Patrick [Dal41]. Paul [Ano34e, Jol51, Rid42]. Pavlov [Anr36]. Pearson [YF36]. Pembrey [Dou35a]. Percy [Car36, F32a, Fea36, Gar48]. Perkin [Rob39]. Perrin [Tow43]. Petavel [Rob36]. Peter [Hin47]. Petrie [Smi45]. Petrovich [Anr36]. Petrovitch [DKY52]. Philemon [Wat40b]. Philip [Ege42, Smi50, Whi49]. Picard [Had42]. Pickard [Ken50]. Pièrre [Ano34e]. Pieter [Str44b]. Pilgrim [FCH44]. Planck [Bor48]. Plaskett [Jon42]. Playfair [Dal41]. Plummer [Gre48]. Pocklington [Ros53]. Pocock [Hin48]. Pope [Gib41]. Porter [Ran40]. Poulton [Car44]. Prain [Bur44a]. Preface [Ano32]. Price [Chr47b, Ken51]. Prince [Ano34a]. Prior [Spe36]. Professor [Bai39, Con39]. Provan [Wis54]. Pyman [Kin44].

Ralph [Mil44a, Mil45a, Mil45b, Whi40]. Ramon [She35]. Ramsay [Gre39b]. Randal [Har42]. Ranken [Gra45, Lai36]. Raper [Har53]. Rawdon [Har42]. Rayleigh [Ege49]. Regan [BN43]. Reginald [B.33a, Bro45, Hin48, Pip52]. Reid [CG48, Kei44]. Rendle [Pra39]. Richard [H.32b, Dav36, Kei54, RS36, Rob53, Str53]. Richardson [Gol54]. Richmond [MW48]. Rickard [Cat35]. Ritter [B.33b]. Robb [Lar38]. Robert [Dam44, JF48, Lin52, Tho39, Des41, Far49, Har41, Ken50, Mun53, Ste40, Wat52]. Roberts [Rid44]. Robertson [Far49]. Robin [Imm38]. Robinson [Har41, Rus51]. Rogers [D.34a, Toi48]. Ronald [N.33]. Rookes [Rus41]. Rose [Ell35]. Rosenhain [D.34b]. Ross [N.33]. Rothschild [Jor38]. Roux [Ano34e]. Rowe [Rod48]. Rowland [Ano50b, Eng50]. Russell [Cam43, Mit41, Whi42]. Rutherford [EC38].

S [Guy41]. Sabatier [Rid42]. Sahni [Ham50]. Sampson [Whi40]. Samuel [Bar48, Ben53, Sut51, Tho49]. Sand [Rus48]. Santiago [She35]. Schack [Hil50]. Schafer [Hil35b]. Schott [Con39]. Schuster [Sim35]. Scott [Dou36, Lid52, Rob48, S.34a, Ste40]. Seligman [Mye41]. Sergei [Tho53b]. Seward [Tho41a]. Seymour [Dou35a]. Sharpey [Hil35b]. Sharpey-Schafer [Hil35b]. Shaw [Gol45, Guy41]. Shearer [Gra42]. Sherrington [Lid52]. Sidgwick [Tiz54]. Sidney [Cal51, Gre51, LC44, R.34]. Sigmund [Tan41]. Simon [Rou49b]. Sinnatt [Ege43]. Sir [And38, Ano50b, B.32a, B32b, B.32d, D’E32, D’E36, D’E40, E.33, Edd40, Ell35, Eve35, GP32, G.34, Gib38, Gib40, Gib41, H.32b, Hil35b, Hol39, Ken35, LG35, Maw35, Moo38, N.33, R.33a, Riv39, Sah40, Sim35, Smi38, T.34, Tho41b, Tho39, Wil38]. Slater [Ken51]. Smiles [Ben53]. Smith [Coo45, Fea38, Gib52, Nut35, Rea48, Tho49, Wil38]. Smithells [Rap40]. Smuts [And52a]. Smyth [Rap41]. Sollas [WW38]. Solomon [Wen38]. Sommerfeld [Bor52]. Sowerby [Bak38]. Spearman [Tho47]. Stanley [Cha50, FC47, Har53, Jon42, Plu45, Sim50, Cal47, Gib52]. Stapf [H.33a]. Steele [M.34b]. Steenberg [Hil50]. Stephen [Bee47]. Stephens [Chr47a]. Stephenson [K.33, Rob49]. Stewart [Fer45, Fos49, Lai36]. Stoney [Dow42]. Stradling [Pip52]. Sturdy [Ege43]. Sugden [Sut51]. Sydney [Atk38, Chr47b, Gar41c, MC48].

T [H32a]. Tannatt [Maw35]. Tate [BN43]. Taylor [Twy38]. Tennyson [Lil51]. Tetley [RS36]. Theobald [Nut35]. Theodore [Mar38]. Thomas [Bak44, Gan53, Gar41b, Har35, Har42, Hil49, Irv41, T.34, Tho41b, DG45, Ell45, Fer48, Fis47, Gra45, Kei48, Ken51, Pic48, Pop38, Rus44, Sum38]. Thompson [Dob49]. Thomson [J.33b, Str41]. Thorpe [Ing41]. Threlfall [H.32b]. Thurland [Spe36]. Tillyard [Imm38]. Toit [Hau49]. Tom [Moo53, Baw54]. Tomlinson [J33a]. Topley [Gre44]. Towle [D.32a]. Trotter [Ell41]. Troup [Ste40]. Tulley [W.32a]. Tullio [Hod42]. Turner [Fos52]. Tutton [Ano39]. Twort [Fil51].

Udny [Yat52]. Unwin [Wil34].

Vavilov [Har54]. Veley [Gar34]. Vernon [Str44a]. Victor [Gar34, Til48, Whi39]. Vilhelm [Gol51]. Vincent [Dal42, Tiz54]. Vines [R.34]. Viscount [Dal42]. Vito [Whi41]. Volterra [Whi41]. Vries [Hal35].

Walcot [Jon43]. Waldemar [Til41]. Walker [Ken35, Rot41]. Wall [Sal52]. Wallace [Dys39, Imm41]. Walter [B.32c, E.33, Tho49, D.34b, Dal47, Hil41, Hir51]. Walther [LS42]. Ward [Dal35]. Warrington [Wen44]. Watson [B.32d, Chr47a, Edd40]. Watts [Bos48]. Waymouth [CG48]. Weber [Tho38]. Wedderburn [Tay49]. Weiss [Tho53a]. Wellcome [Wen38]. Wells [Cal47]. Went [Bri36]. Wentworth [Dob49]. Wenyon [Hoa49]. White [Smi50]. Whitehead [Bos48, Whi48]. Whiteman [Gre44]. Wilfred [Ell41, Tie52]. Wilhelm [Bor52, Tho38]. Willey [Ker43]. William [A.32, Ano34b, B.32d, Bla53, Bro42, Chr47a, Dar40, Fil51, Gib38, Gib41, Gre39a, Gri41, Hev48, Hop53, MB33, Maw35, MW48, Ran40, Ric48, Rid53, Toi48, Whi39, Whi51, Whi54, AL43, And52b, And53, Ben51, Bos48, Con35, Dam44, Des44, Dys39, Fin39, Gan53, Gol45, GS40, Gre44, HS34, Har43, Ing36, Jon38, LD41, Mil35, Pop35, Rod51, Smi45, Tay54, Twy38, W.32b, Wil34, WW38]. Williams [Bla47, Can41]. Willstätter [Rob53]. Wilson [Hil49, Mor40]. Winogradsky [Tho53b]. Wooding [Rus51]. Woods [Bul53]. Woodward [Coo45]. Woollard [Cla40]. Wright [Col48]. Wyndham [Ano50b]. Wynne [Rod51].

Yarrow [D’E32]. Yorke [Wen44]. Young [Atk38, Har43, Tur41]. Yule [Yat52].

Zeeman [Str44b].

References

A:1932:DWD

[A.32] E. V. A. David William Dye, 1887–1932. Obituary Notices of Fellows of the Royal Society, 1(1):75–78, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Andrade:1943:WHB

[AL43] Edward Neville Da Costa Andrade and Kathleen Yardley Lonsdale. William Henry Bragg, 1862–1942. Obituary Notices of Fellows of the Royal Society, 4(12):277–300, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Allen:1947:CGB

[All47] Herbert Stanley Allen. Charles Glover Barkla, 1877–1944. Obituary Notices of Fellows of the Royal Society, 5(15):341–366, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Andrade:1938:SHC

[And38] Edward Neville Da Costa Andrade. Sir Henry Capel Lofft Holden, 1856–1937. Obituary Notices of Fellows of the Royal Society, 2(6):366– 369, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anderson:1952:JCS

[And52a] John Anderson. Jan Christiaan Smuts 1870–1950. Obituary Notices of Fellows of the Royal Society, 8(21):271–273, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Andrewes:1952:WJE

[And52b] Christopher Howard Andrewes. William Joseph Elford 1900–1952. Obituary Notices of Fellows of the Royal Society, 8(21):149–158, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Andrewes:1953:WEG

[And53] Christopher Howard Andrewes. William Ewart Gye, 1884–1952. Obituary Notices of Fellows of the Royal Society, 8(22):418–430, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1932:P

[Ano32] Anonymous. Preface. Obituary Notices of Fellows of the Royal Society, 1(1):2, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1934:APC

[Ano34a] Anonymous. Arthur Prince Chattock, 1860–1934. Obituary Notices of Fellows of the Royal Society, 1(3):292–298, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1934:EWH

[Ano34b] Anonymous. Ernest William Hobson, 1856–1933. Obituary Notices of Fellows of the Royal Society, 1(3):236–249, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1934:JEM

[Ano34c] Anonymous. John Edward Marr, 1857–1933. Obituary Notices of Fellows of the Royal Society, 1(3):250–257, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1934:JJ

[Ano34d] Anonymous. John Joly, 1857–1933. Obituary Notices of Fellows of the Royal Society, 1(3):258–286, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1934:PPE

[Ano34e] Anonymous. Pièrre Paul Émile Roux, 1853–1933. Obituary Notices of Fellows of the Royal Society, 1(3):197–204, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1939:AEH

[Ano39] Anonymous. Alfred Edwin Howard Tutton, 1864–1938. Obituary Notices of Fellows of the Royal Society, 2(7):621–626, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1944:C

[Ano44] Anonymous. Corrigenda. Obituary Notices of Fellows of the Royal Society, 4(13):811, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1949:C

[Ano49] Anonymous. Corrigendum. Obituary Notices of Fellows of the Royal Society, 6(18):661, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1950:CD

[Ano50a] Anonymous. Clifford Dobell 1886–1949. Obituary Notices of Fellows of the Royal Society, 7(19):35–61, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1950:SWR

[Ano50b] Anonymous. Sir Wyndham Rowland Dunstan. Obituary Notices of Fellows of the Royal Society, 7(19):62–81, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anonymous:1952:EHF

[Ano52] Anonymous. Ernest Harold Farmer 1890–1952. Obituary Notices of Fellows of the Royal Society, 8(21):159–169, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Anrep:1936:IPP

[Anr36] Gleb Anrep. Ivan Petrovich Pavlov, 1849–1936. Obituary Notices of Fellows of the Royal Society, 2(5):1–18, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Armstrong:1935:CFC

[Arm35] Edward Frankland Armstrong. Charles Frederick Cross, 1855–1935. Obituary Notices of Fellows of the Royal Society, 1(4):458–464, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Armstrong:1945:MOF

[AS45] Edward Frankland Armstrong and John Lionel Simonsen. Martin Onslow Forster, 1872–1945. Obituary Notices of Fellows of the Royal Society, 5(14):243–261, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Atkins:1938:SY

[Atk38] William Ringrose Gelston Atkins. Sydney Young, 1857–1937. Obituary Notices of Fellows of the Royal Society, 2(6):370–379, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Atkins:1954:HHD

[Atk54] William Ringrose Gelston Atkins. Henry Horatio Dixon, 1869–1953. Obituary Notices of Fellows of the Royal Society, 9(1):78–97, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1932:SDB

[B.32a] J. R. B. Sir David Bruce, 1855–1931. Obituary Notices of Fellows of the Royal Society, 1(1):79–85, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1932:SOJ

[B32b] J. R. B. Sir Otto John Beit, 1865–1930. Obituary Notices of Fellows of the Royal Society, 1(1):60–62, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1932:JWG

[B.32c] P. G. H. B. John Walter Gregory, 1864–1932. Obituary Notices of Fellows of the Royal Society, 1(1):53–59, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1932:SWW

[B.32d] W. B. Sir William Watson Cheyne, Baronet, 1852–1932. Obituary Notices of Fellows of the Royal Society, 1(1):26–30, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1933:HRA

[B.33a] C. V. B. Henry Reginald Arnulph Mallock, 1851–1933. Obituary Notices of Fellows of the Royal Society, 1(2):95–100, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1933:KRG

[B.33b] F. O. B. Karl Ritter von Goebel, 1855–1932. Obituary Notices of Fellows of the Royal Society, 1(2):103–108, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

B:1933:GHB

[B.33c] L. B. George Hartley Bryan, 1864–1928. Obituary Notices of Fellows of the Royal Society, 1(2):139–142, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baddiley:1943:AGG

[Bad43] J. Baddiley. Arthur George Green, 1864–1941. Obituary Notices of Fellows of the Royal Society, 4(12):251–270, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bailey:1939:PAH

[Bai39] Edward Battersby Bailey. Professor Albert Heim, 1849–1937. Obituary Notices of Fellows of the Royal Society, 2(7):471–474, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bailey:1943:GJG

[Bai43] Edward Battersby Bailey. Gerard Jacob De Geer 1858–1943. Obituary Notices of Fellows of the Royal Society, 4(12):475–481, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bailey:1954:ML

[Bai54] Edward Battersby Bailey. Maurice Lugeon 1870–1953. Obituary Notices of Fellows of the Royal Society, 9(1):164–173, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baker:1938:FSM

[Bak38] Henry Frederick Baker. Francis Sowerby Macaulay, 1862–1937. Obituary Notices of Fellows of the Royal Society, 2(6):356–361, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baker:1944:GTB

[Bak44] Henry Frederick Baker. Geoffrey Thomas Bennett 1868–1943. Obituary Notices of Fellows of the Royal Society, 4(13):596–615, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baker:1953:CEI

[Bak53] John Fleetwood Baker. Charles Edward Inglis, 1875–1952. Obituary Notices of Fellows of the Royal Society, 8(22):444–457, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baly:1943:JNC

[Bal43] Edward Charles Cyril Baly. John Norman Collie 1859–1942. Obituary Notices of Fellows of the Royal Society, 4(12):329–356, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Baly:1944:CC

[Bal44] Edward Charles Cyril Baly. Clive Cuthbertson, 1863–1943. Obituary Notices of Fellows of the Royal Society, 4(13):629–654, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bartlett:1948:CSM

[Bar48] Frederic Charles Bartlett. Charles Samuel Myers, 1873–1946. Obituary Notices of Fellows of the Royal Society, 5(16):767–777, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bawden:1954:TG

[Baw54] Frederick Charles Bawden. Tom Goodey 1885–1953. Obituary Notices of Fellows of the Royal Society, 9(1):141–152, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bedson:1947:JCG

[Bed47] Samuel Phillips Bedson. John Charles Grant Ledingham, 1875–1944. Obituary Notices of Fellows of the Royal Society, 5(15):325–340, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Beer:1947:ESG

[Bee47] Gavin Rylands De Beer. Edwin Stephen Goodrich, 1868–1946. Obituary Notices of Fellows of the Royal Society, 5(15):477–490, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Benson:1951:WBB

[Ben51] William Noel Benson. William Blaxland Benham, 1860–1950. Obituary Notices of Fellows of the Royal Society, 7(20):293–301, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bennett:1953:SS

[Ben53] George Macdonald Bennett. Samuel Smiles, 1877–1953. Obituary Notices of Fellows of the Royal Society, 8(22):583–600, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Best:1942:FGB

[Bes42] Charles Herbert Best. Frederick Grant Banting, 1891–1941. Obituary Notices of Fellows of the Royal Society, 4(11):20–26, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Blackman:1945:JBF

[Bla45] Vernon Herbert Blackman. John Bretland Farmer, 1865–1944. Obituary Notices of Fellows of the Royal Society, 5(14):17–31, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Blackett:1947:EJW

[Bla47] P. M. S. Blackett. Evan James Williams, 1903–1945. Obituary Notices of Fellows of the Royal Society, 5(15):386–406, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Blackman:1953:FWK

[Bla53] Vernon Herbert Blackman. Frederick William Keeble, 1870–1952. Obituary Notices of Fellows of the Royal Society, 8(22):490–501, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Burne:1943:CTR

[BN43] Richard Higgins Burne and J. R. Norman. Charles Tate Regan 1878– 1943. Obituary Notices of Fellows of the Royal Society, 4(12):411–426, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Born:1948:MKE

[Bor48] Max Born. Max Karl Ernst Ludwig Planck, 1858–1947. Obituary Notices of Fellows of the Royal Society, 6(17):161–188, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (elec-

tronic). URL http://jstor.org/doi/10.2307/768916;   http://
www.jstor.org/stable/pdfplus/768916.pdf.

Born:1952:AJW

[Bor52] Max Born. Arnold Johannes Wilhelm Sommerfeld 1868–1951. Obituary Notices of Fellows of the Royal Society, 8(21):274–296, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118

(electronic). URL http://www.jstor.org/stable/768813;   http:
//www.jstor.org/stable/pdfplus/768813.pdf.

Boswell:1948:WWW

[Bos48] Percy George Hamnall Boswell. William Whitehead Watts, 1860–1947. Obituary Notices of Fellows of the Royal Society, 6(17):263–279, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bragg:1941:DLE

[Bra41] William Henry Bragg. David Lindsay, Earl of Crawford and Balcarres, 1871–1940. Obituary Notices of Fellows of the Royal Society, 3(9):405– 407, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Briggs:1936:FAF

[Bri36] George Edward Briggs. Friedrich August Ferdinand Christian Went, 1863–1935. Obituary Notices of Fellows of the Royal Society, 2(5):160– 164, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Briggs:1948:FFB

[Bri48] George Edward Briggs. Frederick Frost Blackman, 1866–1947. Obituary Notices of Fellows of the Royal Society, 5(16):651–658, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Brooks:1942:AWH

[Bro42] Frederick Tom Brooks. Arthur William Hill, 1875–1941. Obituary Notices of Fellows of the Royal Society, 4(11):87–100, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Brooks:1945:AHR

[Bro45] Frederick Tom Brooks. Arthur Henry Reginald Buller, 1874–1944. Obituary Notices of Fellows of the Royal Society, 5(14):50–59, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Brown:1948:BAM

[Bro48] George Lindor Brown. Bryan Ausin McSwiney 1894–1947. Obituary Notices of Fellows of the Royal Society, 6(17):146–160, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Bulman:1953:HW

[Bul53] Oliver Meredith Boone Bulman. Henry Woods, 1868–1952. Obituary Notices of Fellows of the Royal Society, 8(22):645–654, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Burkill:1944:DP

[Bur44a] I. H. Burkill. David Prain, 1857–1944. Obituary Notices of Fellows of the Royal Society, 4(13):746–770, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Burkill:1944:HL

[Bur44b] John Charles Burkill. Henri Lebesgue, 1875–1941. Obituary Notices of Fellows of the Royal Society, 4(13):483–490, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

C:1933:GL

[C.33] E. P. C. Graham Lusk, 1866–1932. Obituary Notices of Fellows of the Royal Society, 1(2):143–146, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Calman:1941:EWM

[Cal41] William Thomas Calman. Ernest William Macbride, 1866–1940. Obituary Notices of Fellows of the Royal Society, 3(10):747–759, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Calman:1947:SWK

[Cal47] William Thomas Calman. Stanley Wells Kemp, 1882–1945. Obituary Notices of Fellows of the Royal Society, 5(15):467–476, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Calman:1951:SFH

[Cal51] William Thomas Calman. Sidney Frederic Harmer, 1862–1950. Obituary Notices of Fellows of the Royal Society, 7(20):359–371, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Campbell:1943:AR

[Cam43] Archibald Campbell. Alexander Russell, 1861–1943. Obituary Notices of Fellows of the Royal Society, 4(12):427–428, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Cameron:1951:LGP

[Cam51] Gordon Roy Cameron. Leonard Gregory Parsons, 1879–1950. Obituary Notices of Fellows of the Royal Society, 7(20):453–468, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Cannon:1941:HWC

[Can41] Walter Bradford Cannon. Harvey (Williams) Cushing 1869–1939. Obituary Notices of Fellows of the Royal Society, 3(9):277–290, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Carpenter:1936:PCG

[Car36] Henry Cort Harold Carpenter. Percy Carlyle Gilchrist, 1851–1935. Obituary Notices of Fellows of the Royal Society, 2(5):19–24, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Carpenter:1942:HE

[Car42] G. D. Hale Carpenter. Harry Eltringham, 1873–1941. Obituary Notices of Fellows of the Royal Society, 4(11):113–128, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Carpenter:1944:EBP

[Car44] G. D. Hale Carpenter. Edward Bagnall Poulton 1856–1943. Obituary Notices of Fellows of the Royal Society, 4(13):655–680, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Cathcart:1935:JJR

[Cat35] Edward Provan Cathcart. John James Rickard Macleod, 1876–1935. Obituary Notices of Fellows of the Royal Society, 1(4):584–589, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Cathcart:1948:EWR

[CG48] Edward Provan Cathcart and R. C. Garry. Edward Waymouth Reid, 1862–1948. Obituary Notices of Fellows of the Royal Society, 6(17): 212–218, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Challenger:1950:FSK

[Cha50] F. Challenger. Frederic Stanley Kipping 1863–1949. Obituary Notices of Fellows of the Royal Society, 7(19):182–219, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Christophers:1947:JWW

[Chr47a] Samuel Rickard Christophers. John William Watson Stephens, 1865– 1946. Obituary Notices of Fellows of the Royal Society, 5(15):524–540, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Christophers:1947:SPJ

[Chr47b] Samuel Rickard Christophers. Sydney Price James, 1870–1946. Obituary Notices of Fellows of the Royal Society, 5(15):507–523, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Clark:1940:HHW

[Cla40] Wilfrid Edward Le Gros Clark. Herbert Henry Woollard, 1889–1939. Obituary Notices of Fellows of the Royal Society, 3(8):89–95, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Clemo:1944:FDC

[Cle44] George Roger Clemo. Frederick Daniel Chattaway, 1860–1944. Obituary Notices of Fellows of the Royal Society, 4(13):713–716, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Colebrook:1948:AEW

[Col48] Leonard Colebrook. Almroth Edward Wright, 1861–1947. Obituary Notices of Fellows of the Royal Society, 6(17):297–314, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Conaway:1935:WMO

[Con35] A. W. Conaway. William McFadden Orr, 1866–1934. Obituary Notices of Fellows of the Royal Society, 1(4):559–562, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Conway:1939:PGS

[Con39] Arthur William Conway. Professor G. A. Schott, 1868–1937. Obituary Notices of Fellows of the Royal Society, 2(7):451–454, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Cooper:1945:ASW

[Coo45] C. Forster Cooper. Arthur Smith Woodward, 1864–1944. Obituary Notices of Fellows of the Royal Society, 5(14):79–112, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

D:1932:FAT

[D.32a] H. H. D. Francis Alexander Towle, 1874–1932. Obituary Notices of Fellows of the Royal Society, 1(1):85–87, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

D:1932:EHG

[D.32b] W. C. D. D. Ernest Howard Griffiths, 1851–1932. Obituary Notices of Fellows of the Royal Society, 1(1):15–18, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

D:1934:LJR

[D.34a] A. L. D. Leonard James Rogers, 1862–1933. Obituary Notices of Fellows of the Royal Society, 1(3):299–301, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

D:1934:WR

[D.34b] C. H. D. Walter Rosenhain, 1875–1934. Obituary Notices of Fellows of the Royal Society, 1(3):353–359, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1935:HWD

[Dal35] Henry Hallett Dale. Harold Ward Dudley, 1887–1935. Obituary Notices of Fellows of the Royal Society, 1(4):595–606, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1939:JJA

[Dal39] Henry Hallett Dale. John Jacob Abel, 1857–1938. Obituary Notices of Fellows of the Royal Society, 2(7):577–585, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1940:GB

[Dal40] Henry Hallett Dale. George Barger, 1878–1939. Obituary Notices of Fellows of the Royal Society, 3(8):63–85, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1941:PPL

[Dal41] Henry Hallett Dale. Patrick Playfair Laidlaw, 1881–1940. Obituary Notices of Fellows of the Royal Society, 3(9):427–447, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1942:EVV

[Dal42] Henry Hallett Dale. Edgar Vincent, Viscount D’Abernon, 1857–1942. Obituary Notices of Fellows of the Royal Society, 4(11):83–86, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1944:HGL

[Dal44] Henry Hallett Dale. Henry George Lyons, 1864–1944. Obituary Notices of Fellows of the Royal Society, 4(13):795–809, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1947:WBC

[Dal47] Henry Hallett Dale. Walter Bradford Cannon, 1871–1945. Obituary Notices of Fellows of the Royal Society, 5(15):407–423, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1948:FGH

[Dal48] Henry Hallett Dale. Frederick Gowland Hopkins 1861–1947. Obituary Notices of Fellows of the Royal Society, 6(17):115–145, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dale:1953:CHK

[Dal53] Henry Hallett Dale. Charles Halliley Kellaway, 1889–1952. Obituary Notices of Fellows of the Royal Society, 8(22):502–521, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dampier:1944:WRB

[Dam44] William Cecil Dampier Dampier. William Robert Bousfield 1854– 1943. Obituary Notices of Fellows of the Royal Society, 4(13):570–576, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Darwin:1940:EWB

[Dar40] Charles Galton Darwin. Ernest William Brown, 1866–1938. Obituary Notices of Fellows of the Royal Society, 3(8):18–22, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Davison:1936:RDO

[Dav36] C. Davison. Richard Dixon Oldham, 1858–1936. Obituary Notices of Fellows of the Royal Society, 2(5):111–113, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

deBroglie:1947:CF

[dB47] Louis de Broglie. Charles Fabry, 1867–1945. Obituary Notices of Fellows of the Royal Society, 5(15):445–450, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Margerie:1952:FAA

[De 52] Emmanuel Marie Pierre Martin Jacquin De Margerie. Francois Antoine Alfred Lacroix 1863–1948. Obituary Notices of Fellows of the Royal Society, 8(21):193–205, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

DEyncourt:1932:SAF

[D’E32] Eustace Henry William Tennyson D’Eyncourt. Sir Alfred Fernandez Yarrow, 1842–1932. Obituary Notices of Fellows of the Royal Society, 1(1):7–11, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

DEyncourt:1936:SGH

[D’E36] Eustace Henry William Tennyson D’Eyncourt. Sir George Hadcock, 1861–1936. Obituary Notices of Fellows of the Royal Society, 2(5):140– 143, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

DEyncourt:1940:SHJ

[D’E40] Eustace Henry William Tennyson D’Eyncourt. Sir Henry John Oram, 1858–1939. Obituary Notices of Fellows of the Royal Society, 3(8):155– 158, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Desch:1938:HLC

[Des38] Cecil Henry Desch. Henry Louis Le Chatelier, 1850–1938. Obituary Notices of Fellows of the Royal Society, 2(6):250–259, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Desch:1941:RAH

[Des41] Cecil Henry Desch. Robert Abbott Hadfield, 1858–1940. Obituary Notices of Fellows of the Royal Society, 3(10):647–664, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Desch:1944:WHH

[Des44] Cecil Henry Desch. William Herbert Hatfield 1882–1943. Obituary Notices of Fellows of the Royal Society, 4(13):616–627, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Desch:1945:GDB

[Des45] Cecil Henry Desch. Guy Dunstan Bengough, 1876–1945. Obituary Notices of Fellows of the Royal Society, 5(14):168–178, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Drury:1945:TL

[DG45] Alan Nigel Drury and Ronald Thomson Grant. Thomas Lewis, 1881– 1945. Obituary Notices of Fellows of the Royal Society, 5(14):179–202, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dingle:1941:AF

[Din41] Herbert Dingle. Alfred Fowler 1868–1940. Obituary Notices of Fellows of the Royal Society, 3(9):483–497, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

DeBurghDaly:1952:BPB

[DKY52] Ivan De Burgh Daly, S. A. Komarov, and E. G. Young. Boris Petrovitch Babkin 1877–1950. Obituary Notices of Fellows of the Royal Society, 8(21):12–23, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dobell:1949:DWT

[Dob49] Clifford Dobell. D’Arcy Wentworth Thompson, 1860–1948. Obituary Notices of Fellows of the Royal Society, 6(18):599–617, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Donnan:1942:HF

[Don42] Frederick George Donnan. Herbert Freundlich, 1880–1941. Obituary Notices of Fellows of the Royal Society, 4(11):27–50, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Donnan:1948:ECC

[Don48a] Frederick George Donnan. Edward Charles Cyril Baly 1871–1948. Obituary Notices of Fellows of the Royal Society, 6(17):6–21, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Donnan:1948:EJC

[Don48b] Frederick George Donnan. Ernst Julius Cohen, 1869–1944. Obituary Notices of Fellows of the Royal Society, 5(16):666–687, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Douglas:1935:MSP

[Dou35a] Claude Gordon Douglas. Marcus Seymour Pembrey, 1868–1934. Obituary Notices of Fellows of the Royal Society, 1(4):563–567, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Douglas:1935:GD

[Dou35b] Stewart Ranken Douglas. Georges Dreyer, 1873–1934. Obituary Notices of Fellows of the Royal Society, 1(4):568–576, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Douglas:1936:JSH

[Dou36] Claude Gordon Douglas. John Scott Haldane, 1860–1936. Obituary Notices of Fellows of the Royal Society, 2(5):114–139, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Douglas:1953:LEH

[Dou53] Claude Gordon Douglas. Leonard Erskine Hill, 1866–1952. Obituary Notices of Fellows of the Royal Society, 8(22):431–443, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dowson:1942:GGS

[Dow42] Robert Dowson. George Gerald Stoney, 1863–1942. Obituary Notices of Fellows of the Royal Society, 4(11):183–196, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Dyson:1939:WWC

[Dys39] Frank Watson Dyson. William Wallace Campbell, 1862–1938. Obituary Notices of Fellows of the Royal Society, 2(7):613–619, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

E:1933:SWM

[E.33] T. R. E. Sir Walter Morley Fletcher, 1873–1933. Obituary Notices of Fellows of the Royal Society, 1(2):153–163, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Eve:1938:LR

[EC38] Arthur Stewart (Arthur Stewart) Eve and James Chadwick. Lord Rutherford, 1871–1937. Obituary Notices of Fellows of the Royal Society, 2(6):394–423, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic). URL http://rsbm.

royalsocietypublishing.org/content/royobits/2/6/394;   http:
//www.jstor.org/stable/769080.

Eccles:1945:JAF

[Ecc45] William Henry Eccles. John Ambrose Fleming, 1849–1945. Obituary Notices of Fellows of the Royal Society, 5(14):231–242, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Eddington:1940:SFW

[Edd40] Arthur Stanley Eddington. Sir Frank Watson Dyson, 1868–1939. Obituary Notices of Fellows of the Royal Society, 3(8):159–172, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Eddington:1942:JL

[Edd42] Arthur Stanley Eddington. Joseph Larmor, 1857–1942. Obituary Notices of Fellows of the Royal Society, 4(11):197–207, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Edelsten:1950:JCF

[Ede50] H. Mcd. Edelsten. John Claud Fortescue Fryer 1886–1948. Obituary Notices of Fellows of the Royal Society, 7(19):94–106, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Edgell:1951:AMF

[Edg51] John Augustine Edgell. Arthur Mostyn Field, 1855–1950. Obituary Notices of Fellows of the Royal Society, 7(20):355–358, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Edwards:1941:HCH

[Edw41] Charles Alfred Edwards. Henry Cort Harold Carpenter, 1875–1940. Obituary Notices of Fellows of the Royal Society, 3(10):611–625, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Egerton:1942:JCP

[Ege42] Alfred Charles Egerton. James Charles Philip, 1873–1941. Obituary Notices of Fellows of the Royal Society, 4(11):51–62, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Egerton:1943:FSS

[Ege43] Alfred Charles Egerton. Frank Sturdy Sinnatt 1880–1943. Obituary Notices of Fellows of the Royal Society, 4(12):429–445, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Egerton:1949:LR

[Ege49] Alfred Charles Egerton. Lord Rayleigh, 1875–1947. Obituary Notices of Fellows of the Royal Society, 6(18):502–538, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Elliott:1935:SJR

[Ell35] Thomas Renton Elliott. Sir John Rose Bradfor, Bart. 1863–1935. Obituary Notices of Fellows of the Royal Society, 1(4):527–535, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Elliott:1941:WBL

[Ell41] Thomas Renton Elliott. Wilfred Batten Lewis Trotter, 1872–1939. Obituary Notices of Fellows of the Royal Society, 3(9):325–344, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Elliott:1945:TB

[Ell45] Thomas Renton Elliott. Thomas Barlow, 1845–1945. Obituary Notices of Fellows of the Royal Society, 5(14):158–167, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Elliott:1948:CB

[Ell48] Thomas Renton Elliott. Charles Bolton 1870–1947. Obituary Notices of Fellows of the Royal Society, 6(17):22–36, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Engledow:1950:RHB

[Eng50] F. L. Engledow. Rowland Harry Biffen 1874–1949. Obituary Notices of Fellows of the Royal Society, 7(19):9–25, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Erdelyi:1947:HB

[Erd47] Arthur Erdelyi. Harry Bateman, 1882–1946. Obituary Notices of Fellows of the Royal Society, 5(15):590–618, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Eve:1935:SJC

[Eve35] Arthur Stewart Eve. Sir John Cunningham McLennan, 1867–1935. Obituary Notices of Fellows of the Royal Society, 1(4):577–583, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

F:1932:PG

[F32a] J. B. F. Percy Goom, 1865–1931. Obituary Notices of Fellows of the Royal Society, 1(1):63–64, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

F:1932:AB

[F.32b] J. W. F. Archibald Barr, 1855–1931. Obituary Notices of Fellows of the Royal Society, 1(1):31–36, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Farmer:1949:RR

[Far49] R. C. Farmer. Robert Robertson, 1869–1949. Obituary Notices of Fellows of the Royal Society, 6(18):539–561, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Forster-Cooper:1947:JSG

[FC47] Clive Forster-Cooper. John Stanley Gardiner, 1872–1946. Obituary Notices of Fellows of the Royal Society, 5(15):541–553, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Forster-Cooper:1944:GEP

[FCH44] Clive Forster-Cooper and A. T. Hopwood. Guy Ellcock Pilgrim 1875– 1943. Obituary Notices of Fellows of the Royal Society, 4(13):577–590, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fearnsides:1936:PFK

[Fea36] William George Fearnsides. Percy Fry Kendall, 1856–1936. Obituary Notices of Fellows of the Royal Society, 2(5):57–65, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fearnsides:1938:DBS

[Fea38] William George Fearnsides. Dr. Bernard Smith, 1881–1938. Obituary Notices of Fellows of the Royal Society, 2(6):239–249, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fermor:1945:CSM

[Fer45] Lewis Leigh Fermor. Charles Stewart Middlemiss, 1859–1945. Obituary Notices of Fellows of the Royal Society, 5(14):262–286, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fermor:1948:THH

[Fer48] Lewis Leigh Fermor. Thomas Henry Holland, 1868–1947. Obituary Notices of Fellows of the Royal Society, 6(17):83–114, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fildes:1951:FWT

[Fil51] Paul Gordon Fildes. Frederick William Twort, 1877–1950. Obituary Notices of Fellows of the Royal Society, 7(20):505–517, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Finch:1939:WAB

[Fin39] George Ingle Finch. William Arthur Bone, 1871–1938. Obituary Notices of Fellows of the Royal Society, 2(7):587–611, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Findlay:1951:JAM

[Fin51] G. M. Findlay. James Alexander Murray, 1873–1950. Obituary Notices of Fellows of the Royal Society, 7(20):445–452, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fisher:1947:THM

[Fis47] R. A. Fisher. Thomas Hunt Morgan, 1866–1945. Obituary Notices of Fellows of the Royal Society, 5(15):451–466, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fleure:1941:ACH

[Fle41a] Herbert John Fleure. Alfred Cort Haddon, 1855–1940. Obituary Notices of Fellows of the Royal Society, 3(9):449–465, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fleure:1941:JGF

[Fle41b] Herbert John Fleure. James George Frazer, 1854–1941. Obituary Notices of Fellows of the Royal Society, 3(10):897–914, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Flett:1943:FDA

[Fle43] John Smith Flett. Frank Dawson Adams 1859–1942. Obituary Notices of Fellows of the Royal Society, 4(12):381–393, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Foster:1949:ASE

[Fos49] John Stuart Foster. Arthur Stewart Eve, 1862–1948. Obituary Notices of Fellows of the Royal Society, 6(18):396–407, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (elec-

tronic). URL http://rsbm.royalsocietypublishing.org/content/
6/18/396;   http://www.jstor.org/stable/768932.

Foster:1952:HTB

[Fos52] John Stuart Foster. Howard Turner Barnes 1873–1950. Obituary Notices of Fellows of the Royal Society, 8(21):24–35, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Freeth:1954:AJA

[Fre54] Francis Arthur Freeth. Arthur John Allmand, 1885–1951. Obituary Notices of Fellows of the Royal Society, 9(1):2–13, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Farren:1935:HG

[FT35] William Scott Farren and Henry Thomas Tizard. Herman Glauert, 1892–1934. Obituary Notices of Fellows of the Royal Society, 1(4):607– 610, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Fleming:1938:GF

[FT38] J. A. Fleming and D’Arcy W. Thompson. George Forbes, 1849–1938. Obituary Notices of Fellows of the Royal Society, 2(6):282–286, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

G:1934:SAC

[G.34] M. H. G. Sir Alexander Cruikshank Houston, 1865–1933. Obituary Notices of Fellows of the Royal Society, 1(3):334–344, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gannon:1953:WTC

[Gan53] H. Graham Gannon. William Thomas Calman, 1871–1952. Obituary Notices of Fellows of the Royal Society, 8(22):355–372, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gardner:1934:VHV

[Gar34] J. A. Gardner. Victor Herbert Veley, 1856–1933. Obituary Notices of Fellows of the Royal Society, 1(3):228–235, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gardiner:1941:AGB

[Gar41a] John Stanley Gardiner. Alfred Gibbs Bourne, 1859–1940. Obituary Notices of Fellows of the Royal Society, 3(10):545–549, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gardiner:1941:ATM

[Gar41b] John Stanley Gardiner. Arthur Thomas Masterman, 1869–1941. Obituary Notices of Fellows of the Royal Society, 3(10):813–817, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gardiner:1941:SJH

[Gar41c] John Stanley Gardiner. Sydney John Hickson, 1859–940. Obituary Notices of Fellows of the Royal Society, 3(9):383–394, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Garner:1948:PFF

[Gar48] William Edward Garner. Percy Faraday Frankland, 1858–1946. Obituary Notices of Fellows of the Royal Society, 5(16):697–715, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Garrod:1954:MHG

[Gar54] Lawrence P. Garrod. Mervyn Henry Gordon 1872–1953. Obituary Notices of Fellows of the Royal Society, 9(1):153–163, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gregory:1941:OJL

[GF41] Richard Arman Gregory and Allan Ferguson. Oliver Joseph Lodge, 1851–1940. Obituary Notices of Fellows of the Royal Society, 3(10): 551–574, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gibson:1948:EFA

[GH48] Charles Stanley Gibson and Thomas Percy Hilditch. Edward Frankland Armstrong, 1878–1945. Obituary Notices of Fellows of the Royal Society, 5(16):619–633, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Graaff-Hunter:1953:CFA

[GH53] James De Graaff-Hunter. Charles Frederick Arden-Close, 1865–1952. Obituary Notices of Fellows of the Royal Society, 8(22):327–339, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gibb:1938:SAB

[Gib38] Alexander Gibb. Sir Alexander Blackie William Kennedy, 1847–1928. Obituary Notices of Fellows of the Royal Society, 2(6):212–223, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gibb:1940:SBM

[Gib40] Alexander Gibb. Sir Basil Mott, 1859–1938. Obituary Notices of Fellows of the Royal Society, 3(8):23–38, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gibson:1941:SWJ

[Gib41] Charles Stanley Gibson. Sir William Jackson Pope 1870–1939. Obituary Notices of Fellows of the Royal Society, 3(9):291–324, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gibb:1952:SSC

[Gib52] Claude Dixon Gibb. Stanley Smith Cook 1875–1952. Obituary Notices of Fellows of the Royal Society, 8(21):118–127, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Glazebrook:1935:JAE

[Gla35] Richard Tetley Glazebrook. James Alfred Ewing, 1855–1935. Obituary Notices of Fellows of the Royal Society, 1(4):475–492, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gold:1945:WNS

[Gol45] Ernest Gold. William Napier Shaw, 1854–1945. Obituary Notices of Fellows of the Royal Society, 5(14):203–230, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gold:1951:VFK

[Gol51] Ernest Gold. Vilhelm Friman Koren Bjerknes, 1862–1951. Obituary Notices of Fellows of the Royal Society, 7(20):303–317, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gold:1954:LFR

[Gol54] Ernest Gold. Lewis Fry Richardson, 1881–1953. Obituary Notices of Fellows of the Royal Society, 9(1):216–235, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

G:1932:SFA

[GP32] M. H. G. and E. B. P. Sir Frederick Andrewes, 1859–1932. Obituary Notices of Fellows of the Royal Society, 1(1):37–44, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gray:1942:CS

[Gra42] James Gray. Cresswell Shearer, 1874–1941. Obituary Notices of Fellows of the Royal Society, 4(11):15–19, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Grant:1944:CCF

[Gra44] Kerr Grant. Clinton Coleridge Farr, 1866–1943. Obituary Notices of Fellows of the Royal Society, 4(13):503–506, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Grant:1945:TRL

[Gra45] Kerr Grant. Thomas Ranken Lyle, 1860–1944. Obituary Notices of Fellows of the Royal Society, 5(14):32–49, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Green:1939:JWM

[Gre39a] A. T. Green. Joseph William Mellor, 1869–1938. Obituary Notices of Fellows of the Royal Society, 2(7):573–576, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gregory:1939:JRM

[Gre39b] Richard Arman Gregory. James Ramsay MacDonald, 1866–1937. Obituary Notices of Fellows of the Royal Society, 2(7):475–482, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gregory:1943:EHA

[Gre43] Richard Arman Gregory. Edward Heron-Allen 1861–1943. Obituary Notices of Fellows of the Royal Society, 4(12):447–454, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Greenwood:1944:WWC

[Gre44] Major Greenwood. William Whiteman Carlton Topley, 1886–1944. Obituary Notices of Fellows of the Royal Society, 4(13):698–712, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Greaves:1948:HCP

[Gre48] William Michael Herbert Greaves. Henry Crozier Plummer, 1875–1946. Obituary Notices of Fellows of the Royal Society, 5(16):778–789, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gregory:1951:SGB

[Gre51] Richard Arman Gregory. Sidney George Brown, 1873–1948. Obituary Notices of Fellows of the Royal Society, 7(20):319–327, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Griffiths:1941:GWC

[Gri41] Ezer Griffiths. George William Clarkson Kaye, 1880–1941. Obituary Notices of Fellows of the Royal Society, 3(10):881–895, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Greenwood:1940:WM

[GS40] Major Greenwood and May Smith. William McDougall, 1871–1938. Obituary Notices of Fellows of the Royal Society, 3(8):39–62, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Graham-Smith:1939:GHF

[GSK39] George Stuart Graham-Smith and David Keilin. George Henry Falkiner Nuttall, 1862–1937. Obituary Notices of Fellows of the Royal Society, 2(7):493–499, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Guy:1941:HHS

[Guy41] Henry Lewis Guy. H. S. Hele-Shaw 1854–1941. Obituary Notices of Fellows of the Royal Society, 3(10):791–811, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1932:AB

[H32a] J. P. H. Albert A. T. Brachet, 1869–1930. Obituary Notices of Fellows of the Royal Society, 1(1):64–70, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1932:SRT

[H.32b] W. B. H. Sir Richard Threlfall, 1861–1932. Obituary Notices of Fellows of the Royal Society, 1(1):45–53, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1933:OS

[H.33a] A. W. H. Otto Stapf, 1857–1933. Obituary Notices of Fellows of the Royal Society, 1(2):115–118, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1933:JM

[H.33b] E. W. H. James Mercer, 1883–1932. Obituary Notices of Fellows of the Royal Society, 1(2):164–165, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1933:GCB

[H.33c] S. J. H. Gilbert Charles Bourne, 1861–1933. Obituary Notices of Fellows of the Royal Society, 1(2):126–130, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Haddon:1940:HB

[Had40] Alfred Cort Haddon. Henry Balfour, 1863–1939. Obituary Notices of Fellows of the Royal Society, 3(8):109–115, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hadley:1941:NC

[Had41] W. W. Hadley. Neville Chamberlain, 1869–1940. Obituary Notices of Fellows of the Royal Society, 3(10):731–734, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hadamard:1942:EP

[Had42] Jacques Hadamard. Emile Picard, 1856–1941. Obituary Notices of Fellows of the Royal Society, 4(11):129–150, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hall:1935:HV

[Hal35] A. D. Hall. Hugo de Vries, 1848–1935. Obituary Notices of Fellows of the Royal Society, 1(4):371–373, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hamshaw:1950:BS

[Ham50] H. Thomas Hamshaw. Birbal Sahni, 1891–1949. Obituary Notices of Fellows of the Royal Society, 7(19):264–277, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Harker:1935:HHT

[Har35] Alfred Harker. Herbert Henry Thomas, 1876–1935. Obituary Notices of Fellows of the Royal Society, 1(4):590–594, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Harington:1941:RR

[Har41] Charles Robert Harington. Robert Robinson, 1883–1941. Obituary Notices of Fellows of the Royal Society, 3(10):929–939, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hartley:1942:RTM

[Har42] Harold Brewer Hartley. Randal Thomas Mowbray Rawdon Berkeley, Earl of Berkeley, 1865–1942. Obituary Notices of Fellows of the Royal Society, 4(11):166–182, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hardy:1943:WHY

[Har43] Godfrey Harold Hardy. William Henry Young, 1863–1942. Obituary Notices of Fellows of the Royal Society, 4(12):307–323, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hardy:1950:JH

[Har50] A. C. Hardy. Johan Hjort 1869–1948. Obituary Notices of Fellows of the Royal Society, 7(19):167–181, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hartley:1952:HDD

[Har52] Percival Hartley. Henry Drysdale Dakin 1880–1952. Obituary Notices of Fellows of the Royal Society, 8(21):128–148, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hartley:1953:HSR

[Har53] Percival Hartley. Henry Stanley Raper, 1882–1951. Obituary Notices of Fellows of the Royal Society, 8(22):566–582, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Harland:1954:NIV

[Har54] Sydney Cross Harland. Nicolai Ivanovitch Vavilov, 1885–1942. Obituary Notices of Fellows of the Royal Society, 9(1):259–264, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Haughton:1949:ALD

[Hau49] Sidney Henry Haughton. Alexander Logie Du Toit, 1878–1948. Obituary Notices of Fellows of the Royal Society, 6(18):385–395, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Haworth:1948:JMG

[Haw48] Robert Downs Haworth. John Masson Gulland 1898–1947. Obituary Notices of Fellows of the Royal Society, 6(17):67–82, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hevesy:1948:FWA

[Hev48] George Charles De Hevesy. Francis William Aston, 1877–1945. Obituary Notices of Fellows of the Royal Society, 5(16):634–650, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hickson:1936:JHA

[Hic36] Sydney John Hickson. James Hartley Ashworth, 1874–1936. Obituary Notices of Fellows of the Royal Society, 2(5):204–211, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1935:LC

[Hil35a] A. W. Hill. Leonard Cockayne, 1855–1934. Obituary Notices of Fellows of the Royal Society, 1(4):442–457, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1935:SEA

[Hil35b] Leonard Erskine Hill. Sir Edward Albert Sharpey-Schafer, 1850–1935. Obituary Notices of Fellows of the Royal Society, 1(4):400–407, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1939:EM

[Hil39] Arthur William Hill. Edward Meyrick 1854–1938. Obituary Notices of Fellows of the Royal Society, 2(7):531–548, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1941:WG

[Hil41] Arthur William Hill. Walter Gardiner, 1859–1941. Obituary Notices of Fellows of the Royal Society, 3(10):985–1004, December 1, 1941. CO-DEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hildebrand:1947:GNL

[Hil47] Joel H. Hildebrand. Gilbert Newton Lewis, 1875–1946. Obituary Notices of Fellows of the Royal Society, 5(15):491–506, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1949:JTW

[Hil49] James Peter Hill. James Thomas Wilson, 1861–1945. Obituary Notices of Fellows of the Royal Society, 6(18):643–660, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hill:1950:ASS

[Hil50] A. V. Hill. August Schack Steenberg Krogh, 1874–1949. Obituary Notices of Fellows of the Royal Society, 7(19):220–237, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hindle:1947:PCM

[Hin47] Edward Hindle. Peter Chalmers Mitchell, 1864–1945. Obituary Notices of Fellows of the Royal Society, 5(15):367–372, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hindle:1948:RIP

[Hin48] Edward Hindle. Reginald Innes Pocock 1863–1947. Obituary Notices of Fellows of the Royal Society, 6(17):189–211, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hirst:1951:WNH

[Hir51] Edmund Langley Hirst. Walter Norman Haworth, 1883–1950. Obituary Notices of Fellows of the Royal Society, 7(20):373–404, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hopkins:1942:AH

[HM42] Frederick Gowland Hopkins and Charles James Martin. Arthur Harden, 1865–1940. Obituary Notices of Fellows of the Royal Society, 4(11):2–14, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hoare:1949:CMW

[Hoa49] Cecil Arthur Hoare. Charles Morley Wenyon, 1878–1948. Obituary Notices of Fellows of the Royal Society, 6(18):626–642, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hodge:1942:TLC

[Hod42] William Vallance Douglas Hodge. Tullio Levi-Civita, 1873–1941. Obituary Notices of Fellows of the Royal Society, 4(11):151–165, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hogben:1950:MG

[Hog50] Lancelot Hogben. Major Greenwood 1880–1949. Obituary Notices of Fellows of the Royal Society, 7(19):138–154, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Holmes:1939:SJC

[Hol39] Gordon Morgan Holmes. Sir James Crichton Browne, 1840–1938. Obituary Notices of Fellows of the Royal Society, 2(7):519–521, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Holmes:1941:HH

[Hol41] Gordon Morgan Holmes. Henry Head, 1861–1940. Obituary Notices of Fellows of the Royal Society, 3(10):665–689, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hopkins:1938:AEG

[Hop38] Frederick Gowland Hopkins. Archibald Edward Garrod, 1857–1938. Obituary Notices of Fellows of the Royal Society, 2(6):224–228, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Hopkirk:1953:FWC

[Hop53] K. R. Hopkirk. Frederick William Carter, 1870–1952. Obituary Notices of Fellows of the Royal Society, 8(22):373–388, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

H:1934:WBH

[HS34] F. G. H. and F. E. S. William Bate Hardy, 1864–1933. Obituary Notices of Fellows of the Royal Society, 1(3):326–333, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Holland:1943:HAM

[HS43] Thomas Henry Holland and Leonard James Spencer. Henry Alexander Miers 1858–1942. Obituary Notices of Fellows of the Royal Society, 4(12):369–380, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Imms:1938:RJT

[Imm38] Augustus Daniel Imms. Robin John Tillyard, 1881–1937. Obituary Notices of Fellows of the Royal Society, 2(6):339–345, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Imms:1941:FWE

[Imm41] Augustus Daniel Imms. Frederick Wallace Edwards, 1888–1940. Obituary Notices of Fellows of the Royal Society, 3(10):735–745, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Inglis:1936:WED

[Ing36] Charles Edward Inglis. William Ernest Dalby, 1862–1936. Obituary Notices of Fellows of the Royal Society, 2(5):144–149, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Ingold:1941:JFT

[Ing41] Christopher Kelk Ingold. Jocelyn Field Thorpe, 1872–1940. Obituary Notices of Fellows of the Royal Society, 3(10):531–544, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Irvine:1941:GTM

[Irv41] James Colquhoun Irvine. Gilbert Thomas Morgan, 1872–1940. Obituary Notices of Fellows of the Royal Society, 3(9):355–362, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Irvine:1948:EBB

[Irv48] James Colquhoun Irvine. Earl Baldwin Of Bewdley, K.G. 1867–1947. Obituary Notices of Fellows of the Royal Society, 6(17):2–5, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

J:1933:HT

[J33a] H J. Herbert Tomlinson, 1845–1931. Obituary Notices of Fellows of the Royal Society, 1(2):89–91, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

J:1933:JMT

[J.33b] H. J. John Millar Thomson, 1849–1933. Obituary Notices of Fellows of the Royal Society, 1(2):91–94, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jacob:1947:CB

[Jac47] Charles Jacob. Carles Barrois, 1851–1939. Obituary Notices of Fellows of the Royal Society, 5(15):286–293, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jeffery:1939:LNG

[Jef39] George Barker Jeffery. Louis Napoleon George Filon, 1875–1937. Obituary Notices of Fellows of the Royal Society, 2(7):501–509, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jessop:1953:EGC

[Jes53] H. T. Jessop. Ernest George Coker, 1869–1946. Obituary Notices of Fellows of the Royal Society, 8(22):389–393, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jones:1948:ARH

[JF48] Harold Spencer Jones and Herbert John Fleure. Arthur Robert Hinks, 1873–1945. Obituary Notices of Fellows of the Royal Society, 5(16):716– 732, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Joliot:1951:PL

[Jol51] F. Joliot. Paul Langevin, 1872–1946. Obituary Notices of Fellows of the Royal Society, 7(20):405–419, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jones:1938:WAP

[Jon38] Owen Thomas Jones. William Arthur Parks, 1868–1938. Obituary Notices of Fellows of the Royal Society, 2(6):260–263, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jones:1941:N

[Jon41] Harold Spencer Jones. E. N. Nevill, 1849–1940. Obituary Notices of Fellows of the Royal Society, 3(9):345–353, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jones:1942:JSP

[Jon42] Harold Spencer Jones. John Stanley Plaskett, 1865–1941. Obituary Notices of Fellows of the Royal Society, 4(11):67–82, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jones:1943:WG

[Jon43] Owen Thomas Jones. Walcot Gibson 1864–1941. Obituary Notices of Fellows of the Royal Society, 4(12):271–275, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Jordan:1938:LR

[Jor38] Heinrich Ernst Karl Jordan. Lord Rothschild, 1868–1937. Obituary Notices of Fellows of the Royal Society, 2(6):385–386, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

K:1933:JS

[K.33] S. W. K. John Stephenson, 1871–1933. Obituary Notices of Fellows of the Royal Society, 1(2):149–152, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keeble:1941:HEA

[Kee41] Frederick William Keeble. Henry Edward Armstrong, 1848–1937. Obituary Notices of Fellows of the Royal Society, 3(9):229–245, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keith:1938:JAM

[Kei38] Arthur Keith. John Alexander MacWilliam, 1857–1937. Obituary Notices of Fellows of the Royal Society, 2(6):334–338, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keith:1943:EF

[Kei43] Arthur Keith. Edward Fawcett 1867–1942. Obituary Notices of Fellows of the Royal Society, 4(12):325–328, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keith:1944:JRM

[Kei44] Arthur Keith. James Reid Moir, 1879–1944. Obituary Notices of Fellows of the Royal Society, 4(13):733–745, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keith:1948:THB

[Kei48] Arthur Keith. Thomas Hastie Bryce, 1862–1946. Obituary Notices of Fellows of the Royal Society, 5(16):659–665, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Keith:1954:RHB

[Kei54] Arthur Keith. Richard Higgins Burne, 1868–1953. Obituary Notices of Fellows of the Royal Society, 9(1):26–33, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Kendal:1935:SJW

[Ken35] J. Kendal. Sir James Walker, 1863–1935. Obituary Notices of Fellows of the Royal Society, 1(4):536–549, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Kenyon:1950:RHP

[Ken50] Joseph Kenyon. Robert Howson Pickard, 1874–1949. Obituary Notices of Fellows of the Royal Society, 7(19):252–263, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Kendall:1951:TSP

[Ken51] James Pickering Kendall. Thomas Slater Price, 1875–1949. Obituary Notices of Fellows of the Royal Society, 7(20):469–474, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Kerr:1943:AW

[Ker43] John Graham Kerr. Arthur Willey, 1867–1942. Obituary Notices of Fellows of the Royal Society, 4(12):395–410, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Kemp:1943:EJA

[KH43] Stanley Wells Kemp and Archibald Vivian Hill. Edgar Johnson Allen 1866–1942. Obituary Notices of Fellows of the Royal Society, 4(12): 357–367, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

King:1944:FLP

[Kin44] Harold King. Frank Lee Pyman, 1882–1944. Obituary Notices of Fellows of the Royal Society, 4(13):681–697, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

King:1950:EJG

[Kin50] W. B. R. King. Edmund Johnston Garwood: 1864–1949. Obituary Notices of Fellows of the Royal Society, 7(19):107–113, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

L:1934:ABM

[L.34a] J. B. L. Archibald Byron Macallum, 1858–1934. Obituary Notices of Fellows of the Royal Society, 1(3):287–291, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

L:1934:FAB

[L.34b] W. D. L. Francis Arthur Bather, 1863–1934. Obituary Notices of Fellows of the Royal Society, 1(3):302–314, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Laidlaw:1936:SRD

[Lai36] P. Laidlaw. Stewart Ranken Douglas, 1871–1936. Obituary Notices of Fellows of the Royal Society, 2(5):175–182, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lang:1947:JEH

[Lan47] William Henry Lang. John Ernest Holloway, 1881–1945. Obituary Notices of Fellows of the Royal Society, 5(15):424–444, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lang:1949:FOB

[Lan49] W. H. Lang. Fredrick Orpen Bower, 1855–1948. Obituary Notices of Fellows of the Royal Society, 6(18):346–374, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Larmor:1938:AAR

[Lar38] Joseph Larmor. Alfred Arthur Robb, 1873–1938. Obituary Notices of Fellows of the Royal Society, 2(6):315–321, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lenox-Conyngham:1944:SGB

[LC44] Gerald Ponsonby Lenox-Conyngham. Sidney Gerald Burrard 1860– 1943. Obituary Notices of Fellows of the Royal Society, 4(13):507–522, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Ledingham:1941:WB

[LD41] John Charles Grant Ledingham and Cecil Clifford Dobell. William Bulloch, 1868–1941. Obituary Notices of Fellows of the Royal Society, 3(10):819–852, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Leathes:1940:JM

[Lea40] John Beresford Leathes. John Mellanby, 1878–1939. Obituary Notices of Fellows of the Royal Society, 3(8):173–195, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lewis:1950:GHH

[Lew50] W. B. Lewis. George Hugh Henderson 1892–1949. Obituary Notices of Fellows of the Royal Society, 7(19):155–166, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Love:1935:SHL

[LG35] Augustus Edward Hough Love and Richard Tetley Glazebrook. Sir Horace Lamb, 1849–1934. Obituary Notices of Fellows of the Royal Society, 1(4):374–392, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Liddell:1952:CSS

[Lid52] Edward George Tandy Liddell. Charles Scott Sherrington 1857–1952. Obituary Notices of Fellows of the Royal Society, 8(21):241–270, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lillicrap:1951:ETE

[Lil51] C. S. Lillicrap. Eustace Tennyson D’Eyncourt, 1868–1951. Obituary Notices of Fellows of the Royal Society, 7(20):341–354, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Linstead:1952:GAR

[Lin52] Reginald Patrick Linstead. George Armand Robert Kon 1892–1951. Obituary Notices of Fellows of the Royal Society, 8(21):170–192, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Lindemann:1942:WN

[LS42] Frederick Alexander Lindemann and Franz Eugen Simon. Walther Nernst, 1864–1941. Obituary Notices of Fellows of the Royal Society, 4(11):101–112, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

M:1934:LCA

[M.34a] C. J. M. Léon Charles Albert Calmette, 1863–1933. Obituary Notices of Fellows of the Royal Society, 1(3):315–325, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

M:1934:BDS

[M.34b] D. O. M. Bertram Dillon Steele, 1870–1934. Obituary Notices of Fellows of the Royal Society, 1(3):345–348, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

M:1934:JCE

[M.34c] F. H. A. M. James Cossar Ewart, 1851–1933. Obituary Notices of Fellows of the Royal Society, 1(3):189–195, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Marshall:1938:JTC

[Mar38] C. R. Marshall. John Theodore Cash, 1854–1938. Obituary Notices of Fellows of the Royal Society, 2(6):294–300, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Martin:1939:AEB

[Mar39] Charles James Martin. Arthur Edwin Boycott, 1877–1938. Obituary Notices of Fellows of the Royal Society, 2(7):561–571, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Martin:1945:JAA

[Mar45] Charles James Martin. Joseph Arthur Arkwright, 1864–1944. Obituary Notices of Fellows of the Royal Society, 5(14):126–140, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Mason:1943:EJB

[Mas43] E. W. Mason. Edwin John Butler 1874–1943. Obituary Notices of Fellows of the Royal Society, 4(12):455–474, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Massey:1952:LJC

[Mas52] Harrie Stewart Wilson Massey. Leslie John Comrie 1893–1950. Obituary Notices of Fellows of the Royal Society, 8(21):96–107, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Mawson:1935:STW

[Maw35] Douglas Mawson. Sir Tannatt William Edgeworth David 1858–1934. Obituary Notices of Fellows of the Royal Society, 1(4):493–501, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

M-B:1933:AWA

[MB33] P. M-B. Alfred William Alcock, 1859–1933. Obituary Notices of Fellows of the Royal Society, 1(2):119–126, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

MacNalty:1948:SAM

[MC48] A. S. MacNalty and James Craigie. Sydney Arthur Monckton Copeman 1862–1947. Obituary Notices of Fellows of the Royal Society, 6(17): 37–50, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

McCrea:1951:EAM

[McC51] William Hunter McCrea. Edward Arthur Milne, 1896–1950. Obituary Notices of Fellows of the Royal Society, 7(20):421–443, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Melville:1953:MGE

[Mel53] Harry Work Melville. Meredith Gwynne Evans, 1904–1952. Obituary Notices of Fellows of the Royal Society, 8(22):394–409, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Michell:1941:JHM

[Mic41] Anthony George Maldon Michell. John Henry Michell, 1863–1940. Obituary Notices of Fellows of the Royal Society, 3(9):363–382, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Milner:1935:WMH

[Mil35] Samuel Roslington Milner. William Mitchinson Hicks, 1850–1934. Obituary Notices of Fellows of the Royal Society, 1(4):393–399, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Milne:1941:AEH

[Mil41] Edward Arthur Milne. Augustus Edward Hough Love, 1863–1940. Obituary Notices of Fellows of the Royal Society, 3(9):467–482, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Milne:1944:RHF

[Mil44a] E. A. Milne. Ralph Howard Fowler. Journal of the London Mathematical Society, 19(76 (Part 4)):244–256, October 1944. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic). Reprinted, with minor changes, in [Mil44a]. See also extended version with publication list in [Mil45a].

Milne:1944:HFN

[Mil44b] Edward Arthur Milne. Hugh Frank Newall, 1857–1944. Obituary Notices of Fellows of the Royal Society, 4(13):717–732, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Milne:1945:ONR

[Mil45a] E. A. Milne. Obituary notices: Ralph Howard Fowler [17 January 1889–28 July 1944]. Monthly Notices of the Royal Astronomical Society, 105(2):80–87, April 1945. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic). See also abbreviated versions [Mil44a, Mil45b].

Milne:1945:RHF

[Mil45b] Edward Arthur Milne. Ralph Howard Fowler, 1889–1944. Obituary Notices of Fellows of the Royal Society, 5(14):60–78, November 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic). URL

http://www.jstor.org/stable/769110. Minor changes from [Mil44a].

See also extended version with publication list in [Mil45a]. See also extended version with publication list in [Mil45a].

Milne:1947:JHJ

[Mil47] Edward Arthur Milne. James Hopwood Jeans, 1877–1946. Obituary Notices of Fellows of the Royal Society, 5(15):573–589, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Mitchell:1941:HAR

[Mit41] Peter Chalmers Mitchell. Herbrand Arthur Russell Duke of Bedford, 1858–1940. Obituary Notices of Fellows of the Royal Society, 3(9):499– 502, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Moore:1938:SHJ

[Moo38] H. Moore. Sir Herbert Jackson, 1863–1938. Obituary Notices of Fellows of the Royal Society, 2(6):306–314, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Moore:1953:FTB

[Moo53] W. C. Moore. Frederick Tom Brooks, 1882–1952. Obituary Notices of Fellows of the Royal Society, 8(22):340–354, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Morgan:1940:EBW

[Mor40] Thomas Hunt Morgan. Edmund Beecher Wilson, 1856–1939. Obituary Notices of Fellows of the Royal Society, 3(8):123–138, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Munro:1953:RN

[Mun53] J. W. Munro. Robert Newstead, 1859–1947. Obituary Notices of Fellows of the Royal Society, 8(22):548–553, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Murray:1938:EHK

[Mur38] James Alexander Murray. Edgar Hartley Kettle, 1882–1938. Obituary Notices of Fellows of the Royal Society, 2(6):301–305, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Murray:1941:HAL

[Mur41] G. Murray. Herbert Albert Laurens Fisher, 1865–1940. Obituary Notices of Fellows of the Royal Society, 3(10):519–529, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Murray:1950:JEB

[Mur50] J. A. Murray. Joseph Edwin Barnard 1870–1949. Obituary Notices of Fellows of the Royal Society, 7(19):2–8, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Milne:1948:HWR

[MW48] Edward Arthur Milne and F. Puryer White. Herbert William Richmond, 1863–1948. Obituary Notices of Fellows of the Royal Society, 6(17):219–230, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Myers:1941:CGS

[Mye41] Charles Samuel Myers. Charles Gabriel Seligman, 1873–1940. Obituary Notices of Fellows of the Royal Society, 3(10):627–646, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Myres:1941:AJE

[Myr41] J. L. Myres. Arthur John Evans, 1851–1941. Obituary Notices of Fellows of the Royal Society, 3(10):941–968, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

N:1932:AAM

[N.32] H. F. N. Albert Abraham Michelson, 1852–1931. Obituary Notices of Fellows of the Royal Society, 1(1):18–25, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

N:1933:SRR

[N.33] G. H. F. N. Sir Ronald Ross, 1857–1933. Obituary Notices of Fellows of the Royal Society, 1(2):108–115, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Newall:1939:GEH

[New39] Hugh Frank Newall. George Ellery Hale, 1868–1938. Obituary Notices of Fellows of the Royal Society, 2(7):523–529, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Nuttall:1935:TS

[Nut35] George Henry Falkiner Nuttall. Theobald Smith, 1859–1934. Obituary Notices of Fellows of the Royal Society, 1(4):514–521, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Parsons:1936:CLM

[Par36] John Herbert Parsons. Conwy Lloyd Morgan, 1852–1936. Obituary Notices of Fellows of the Royal Society, 2(5):25–27, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Parkes:1950:FHA

[Par50] Alan Sterling Parkes. Francis Hugh Adam Marshall, 1878–1949. Obituary Notices of Fellows of the Royal Society, 7(19):238–251, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Peters:1954:OM

[Pet54] Rudolph Albert Peters. Otto Meyerhof, 1884–1951. Obituary Notices of Fellows of the Royal Society, 9(1):174–200, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Picken:1948:THL

[Pic48] D. K. Picken. Thomas Howell Laby, 1880–1946. Obituary Notices of Fellows of the Royal Society, 5(16):733–755, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Pippard:1952:RES

[Pip52] A. J. S. Pippard. Reginald Edward Stradling, 1891–1952. Obituary Notices of Fellows of the Royal Society, 8(21):297–307, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Plummer:1945:ASE

[Plu45] Henry Crozier Keating Plummer. Arthur Stanley Eddington, 1882– 1944. Obituary Notices of Fellows of the Royal Society, 5(14):113–125, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Pringsheim:1949:LJ

[PM49] E. G. Pringsheim and Ernest John Maskell. Ludwig Jost, 1865–1947. Obituary Notices of Fellows of the Royal Society, 6(18):471–478, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Pope:1935:WB

[Pop35] W. J. Pope. William Barlow, 1845–1934. Obituary Notices of Fellows of the Royal Society, 1(4):367–370, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Pope:1938:TML

[Pop38] William Jackson Pope. Thomas Martin Lowry, 1874–1938. Obituary Notices of Fellows of the Royal Society, 2(6):287–293, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Poulton:1935:FAD

[Pou35] Edward Bagnall Poulton. Frederick Augustus Dixey, 1855–1935. Obituary Notices of Fellows of the Royal Society, 1(4):465–474, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Prain:1939:ABR

[Pra39] David Prain. Alfred Barton Rendle, 1865–1938. Obituary Notices of Fellows of the Royal Society, 2(7):511–517, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

R:1932:GCD

[R.32] A. B. R. George Claridge Druce, 1850–1932. Obituary Notices of Fellows of the Royal Society, 1(1):12–14, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

R:1933:SDC

[R.33a] H. R. R. Sir Dugald Clerk, 1854–1932. Obituary Notices of Fellows of the Royal Society, 1(2):101–102, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

R:1933:ACC

[R.33b] R. R. Alfred Chaston Chapman, 1869–1932. Obituary Notices of Fellows of the Royal Society, 1(2):146–148, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

R:1934:SHV

[R.34] A. B. R. Sidney Howard Vines, 1849–1934. Obituary Notices of Fellows of the Royal Society, 1(3):185–188, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rankine:1940:AWP

[Ran40] Alexander Oliver Rankine. Alfred William Porter, 1863–1939. Obituary Notices of Fellows of the Royal Society, 3(8):87–88, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Raper:1935:JBC

[Rap35] H. S. Raper. Julius Berend Cohen, 1859–1935. Obituary Notices of Fellows of the Royal Society, 1(4):502–513, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Raper:1940:AS

[Rap40] H. S. Raper. Arthur Smithells, 1860–1939. Obituary Notices of Fellows of the Royal Society, 3(8):97–107, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Raper:1941:JSM

[Rap41] Henry Stanley Raper. John Smyth Macdonald, 1867–1941. Obituary Notices of Fellows of the Royal Society, 3(10):854–866, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Read:1948:JSF

[Rea48] Herbert Harold Read. John Smith Flett, 1869–1947. Obituary Notices of Fellows of the Royal Society, 5(16):688–696, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Read:1952:AM

[Rea52] John Read. Alexander McKenzie 1869–1951. Obituary Notices of Fellows of the Royal Society, 8(21):206–228, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Read:1953:JCI

[Rea53] John Read. James Colquhoun Irvine, 1877–1952. Obituary Notices of Fellows of the Royal Society, 8(22):458–489, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Ricardo:1948:FWL

[Ric48] Harry Ralph Ricardo. Frederick William Lanchester, 1868–1946. Obituary Notices of Fellows of the Royal Society, 5(16):756–766, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rideal:1942:PS

[Rid42] Eric Keightley Rideal. Paul Sabatier, 1859–1941. Obituary Notices of Fellows of the Royal Society, 4(11):63–66, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rideal:1944:JKR

[Rid44] Eric Keightley Rideal. John Keith Roberts, 1897–1944. Obituary Notices of Fellows of the Royal Society, 4(13):789–794, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rideal:1953:JWM

[Rid53] Eric Keightley Rideal. James William McBain, 1882–1953. Obituary Notices of Fellows of the Royal Society, 8(22):529–547, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rivett:1939:SDO

[Riv39] Albert Cherbury David Rivett. Sir David Orme Masson, 1858–1937. Obituary Notices of Fellows of the Royal Society, 2(7):455–464, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robertson:1936:JEP

[Rob36] Robert Robertson. Joseph Ernest Petavel, 1873–1936. Obituary Notices of Fellows of the Royal Society, 2(5):183–203, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robinson:1939:AGP

[Rob39] Robert Robinson. Arthur George Perkin, 1861–1937. Obituary Notices of Fellows of the Royal Society, 2(7):445–450, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robertson:1945:JJF

[Rob45] Robert Robertson. John Jacob Fox, 1874–1944. Obituary Notices of Fellows of the Royal Society, 5(14):141–157, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robinson:1947:AL

[Rob47] Robert Robinson. Arthur Lapworth, 1872–1941. Obituary Notices of Fellows of the Royal Society, 5(15):554–572, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robertson:1948:AS

[Rob48] Robert Robertson. Alexander Scott, 1853–1947. Obituary Notices of Fellows of the Royal Society, 6(17):251–262, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robertson:1949:MS

[Rob49] Muriel Robertson. Marjory Stephenson, 1885–1948. Obituary Notices of Fellows of the Royal Society, 6(18):562–577, November 1, 1949. CO-DEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robertson:1952:GC

[Rob52] A. Robertson. Gilbert Cook 1885–1951. Obituary Notices of Fellows of the Royal Society, 8(21):108–117, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Robinson:1953:RW

[Rob53] R. Robinson. Richard Willstätter, 1872–1942. Obituary Notices of Fellows of the Royal Society, 8(22):609–634, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rodd:1948:FMR

[Rod48] E. H. Rodd. Frederick Maurice Rowe, 1891–1946. Obituary Notices of Fellows of the Royal Society, 6(17):231–250, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rodd:1951:WPW

[Rod51] E. H. Rodd. William Palmer Wynne, 1861–1950. Obituary Notices of Fellows of the Royal Society, 7(20):519–531, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rosenhead:1953:HCP

[Ros53] Louis Rosenhead. Henry Cabourn Pocklington, 1870–1952. Obituary Notices of Fellows of the Royal Society, 8(22):555–565, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rottenburg:1941:MW

[Rot41] H. Rottenburg. Miles Walker, 1868–1941. Obituary Notices of Fellows of the Royal Society, 3(10):779–789, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rous:1947:KL

[Rou47] Francis Peyton Rous. Karl Landsteiner, 1868–1943. Obituary Notices of Fellows of the Royal Society, 5(15):294–324, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Roughton:1949:JB

[Rou49a] Francis John Worsley Roughton. Joseph Barcroft, 1872–1947. Obituary Notices of Fellows of the Royal Society, 6(18):315–345, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rous:1949:SF

[Rou49b] Francis Peyton Rous. Simon Flexner, 1863–1946. Obituary Notices of Fellows of the Royal Society, 6(18):408–445, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Rayleigh:1936:RTG

[RS36] Rayleigh and F. J. Selby. Richard Tetley Glazebrook, 1854–1935. Obituary Notices of Fellows of the Royal Society, 2(5):28–56, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1941:REB

[Rus41] Alexander Russell. Rookes Evelyn Bell Crompton, 1845–1940. Obituary Notices of Fellows of the Royal Society, 3(9):395–403, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1942:ADH

[Rus42] Edward John Russell. Alfred Daniel Hall, 1864–1942. Obituary Notices of Fellows of the Royal Society, 4(11):229–250, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1944:THM

[Rus44] E. J. Russell. Thomas Hudson Middleton 1863–1943. Obituary Notices of Fellows of the Royal Society, 4(13):554–569, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1948:AS

[Rus48] Frederick Stratten Russell. Alexander Sand, 1901–1945. Obituary Notices of Fellows of the Royal Society, 5(16):803–815, May 1, 1948. CO-DEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1951:GWR

[Rus51] Edward John Russell. Gilbert Wooding Robinson, 1888–1950. Obituary Notices of Fellows of the Royal Society, 7(20):475–491, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Russell:1954:JHO

[Rus54] Frederick Stratten Russell. James Herbert Orton, 1884–1953. Obituary Notices of Fellows of the Royal Society, 9(1):201–215, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Ryde:1949:CCP

[Ryd49] John Walter Ryde. Clifford Copland Paterson, 1879–1948. Obituary Notices of Fellows of the Royal Society, 6(18):479–501, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

S:1934:DHS

[S.34a] A. C. S. Dukinfield Henry Scott, 1854–1934. Obituary Notices of Fellows of the Royal Society, 1(3):205–227, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

S:1934:DB

[S.34b] G. E. S. Davidson Black, 1884–1934. Obituary Notices of Fellows of the Royal Society, 1(3):360–365, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Saha:1940:SJC

[Sah40] Meghnad N. Saha. Sir Jagadis Chunder Bose, 1858–1937. Obituary Notices of Fellows of the Royal Society, 3(8):3–12, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Salisbury:1952:FWO

[Sal52] E. J. Salisbury. Francis Wall Oliver 1864–1951. Obituary Notices of Fellows of the Royal Society, 8(21):229–240, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Salisbury:1954:FEF

[Sal54] E. J. Salisbury. Felix Eugene Fritsch 1879–1954. Obituary Notices of Fellows of the Royal Society, 9(1):130–140, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Schonland:1950:HJV

[Sch50] Basil Ferdinand Jamieson Schonland. Hendrik Johannes Van Der Bijl 1887–1948. Obituary Notices of Fellows of the Royal Society, 7(19): 26–34, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Sherrington:1935:SRC

[She35] Charles Scott Sherrington. Santiago Ramon y Cajal, 1852–1934. Obituary Notices of Fellows of the Royal Society, 1(4):424–441, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Simpson:1935:SAS

[Sim35] George Clarke Simpson. Sir Arthur Schuster, 1851–1934. Obituary Notices of Fellows of the Royal Society, 1(4):408–423, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Simonsen:1944:GGH

[Sim44] J. L. Simonsen. George Gerald Henderson 1862–1942. Obituary Notices of Fellows of the Royal Society, 4(13):491–502, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Simonsen:1950:CSG

[Sim50] J. L. Simonsen. Charles Stanley Gibson 1884–1950. Obituary Notices of Fellows of the Royal Society, 7(19):114–137, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Smith:1938:SHE

[Smi38] William Wright Smith. Sir Herbert Eustace Maxwell, 1845–1937. Obituary Notices of Fellows of the Royal Society, 2(6):387–393, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Smith:1939:AH

[Smi39] W. Campbell Smith. Arthur Hutchinson, 1866–1937. Obituary Notices of Fellows of the Royal Society, 2(7):483–491, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Smith:1941:JCB

[Smi41] Frank Edward Smith. John Cadman, Baron Cadman, 1877–1941. Obituary Notices of Fellows of the Royal Society, 3(10):915–927, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Smith:1945:WMF

[Smi45] Sidney Smith. William Matthew Flinders Petrie, 1853–1942. Obituary Notices of Fellows of the Royal Society, 5(14):3–16, November 1, 1945. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Smith:1950:PBW

[Smi50] Wilson Smith. Philip Bruce White, 1891–1949. Obituary Notices of Fellows of the Royal Society, 7(19):278–292, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Soddy:1939:JEM

[Sod39] Frederick Soddy. James Ernest Marsh, 1860–1938. Obituary Notices of Fellows of the Royal Society, 2(7):549–556, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Southwell:1941:CFJ

[Sou41] Richard Vynne Southwell. Charles Frewen Jenkin, 1865–1940. Obituary Notices of Fellows of the Royal Society, 3(10):575–585, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Spencer:1936:GTP

[Spe36] Leonard James Spencer. George Thurland Prior, 1862–1936. Obituary Notices of Fellows of the Royal Society, 2(5):150–159, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Seward:1940:AH

[ST40] Albert Charles Seward and Cecil Edgar Tilley. Alfred Harker, 1859– 1939. Obituary Notices of Fellows of the Royal Society, 3(8):197–216, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Stebbing:1940:RST

[Ste40] E. P. Stebbing. Robert Scott Troup, 1874–1939. Obituary Notices of Fellows of the Royal Society, 3(8):217–219, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Stiles:1939:AJE

[Sti39] Walter Stiles. Alfred James Ewart, 1872–1937. Obituary Notices of Fellows of the Royal Society, 2(7):465–469, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Strutt:1941:JJT

[Str41] Robert John Strutt. Joseph John Thomson, 1856–1940. Obituary Notices of Fellows of the Royal Society, 3(10):587–609, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Strutt:1944:CVB

[Str44a] John William Strutt. Charles Vernon Boys, 1855–1944. Obituary Notices of Fellows of the Royal Society, 4(13):771–788, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Strutt:1944:PZ

[Str44b] Robert John Strutt. Pieter Zeeman 1865–1943. Obituary Notices of Fellows of the Royal Society, 4(13):591–595, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Stratton:1953:RAG

[Str53] Frederick John Marrian Stratton. Richard Arman Gregory, 1864–1952. Obituary Notices of Fellows of the Royal Society, 8(22):410–417, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Stratton:1954:HAD

[Str54] Frederick John Marrian Stratton. Henri Alexandre Deslandres, 1853– 1948. Obituary Notices of Fellows of the Royal Society, 9(1):64–77, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Sumpner:1938:TM

[Sum38] W. E. Sumpner. Thomas Mather, 1856–1937. Obituary Notices of Fellows of the Royal Society, 2(6):380–384, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Sutton:1951:SS

[Sut51] Leslie Ernest Sutton. Samuel Sugden, 1892–1950. Obituary Notices of Fellows of the Royal Society, 7(20):493–503, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Synge:1933:JCF

[Syn33] John Lighton Synge. John Charles Fields, 1863–1932. Obituary Notices of Fellows of the Royal Society, 1(2):131–138, December 1, 1933. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

T:1934:STM

[T.34] H. W. T. Sir Thomas Muir, 1844–1934. Obituary Notices of Fellows of the Royal Society, 1(3):179–184, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tansley:1939:AHC

[Tan39] Arthur George Tansley. Arthur Harry Church, 1865–1937. Obituary Notices of Fellows of the Royal Society, 2(7):432–443, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tansley:1941:SF

[Tan41] Arthur George Tansley. Sigmund Freud 1856–1939. Obituary Notices of Fellows of the Royal Society, 3(9):247–275, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Taylor:1949:JHM

[Tay49] Hugh Stott Taylor. Joseph Henry Maclagen Wedderburn, 1882–1948. Obituary Notices of Fellows of the Royal Society, 6(18):618–625, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Taylor:1954:WCD

[Tay54] Geoffrey Ingram Taylor. William Cecil Dampier, 1867–1952. Obituary Notices of Fellows of the Royal Society, 9(1):54–63, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thorpe:1935:HBB

[Tho35] J. H. Thorpe. Herbert Brereton Baker, 1862–1935. Obituary Notices of Fellows of the Royal Society, 1(4):522–526, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thompson:1938:MWC

[Tho38] D’arcy Wentworth Thompson. Max Wilhelm Carl Weber, 1852–1937. Obituary Notices of Fellows of the Royal Society, 2(6):346–355, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thorpe:1939:SRM

[Tho39] Jocelyn Field Thorpe. Sir Robert Mond, 1867–1938. Obituary Notices of Fellows of the Royal Society, 2(7):627–632, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thomas:1941:ACS

[Tho41a] Hugh Hamshaw Thomas. Albert Charles Seward, 1863–1941. Obituary Notices of Fellows of the Royal Society, 3(10):867–880, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thompson:1941:STL

[Tho41b] D’arcy Wentworth Thompson. Sir Thomas Little Heath 1861–1940. Obituary Notices of Fellows of the Royal Society, 3(9):409–426, January 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thomson:1947:CS

[Tho47] Godfrey Thomson. Charles Spearman, 1863–1945. Obituary Notices of Fellows of the Royal Society, 5(15):373–385, February 1, 1947. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thomas:1949:SWJ

[Tho49] J. S. G. Thomas. Samuel Walter Johnson Smith, 1871–1948. Obituary Notices of Fellows of the Royal Society, 6(18):578–598, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thomas:1953:FEW

[Tho53a] Hugh Hamshaw Thomas. Frederick Ernest Weiss, 1865–1953. Obituary Notices of Fellows of the Royal Society, 8(22):601–608, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Thornton:1953:SNW

[Tho53b] Henry Gerard Thornton. Sergei Nicholaevitch Winogradsky, 1856– 1953. Obituary Notices of Fellows of the Royal Society, 8(22):635–644, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tiegs:1952:WEA

[Tie52] Oscar Werner Tiegs. Wilfred Eade Agar 1882–1951. Obituary Notices of Fellows of the Royal Society, 8(21):2–11, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tilley:1941:WCB

[Til41] Cecil Edgar Tilley. Waldemar Christopher Brögger, 1851–1940. Obituary Notices of Fellows of the Royal Society, 3(10):503–517, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tilley:1948:VMG

[Til48] Cecil Edgar Tilley. Victor Moritz Goldschmidt 1888–1947. Obituary Notices of Fellows of the Royal Society, 6(17):51–66, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Titchmarsh:1949:GHH

[Tit49] E. C. Titchmarsh. Godfrey Harold Hardy, 1877–1947. Obituary Notices of Fellows of the Royal Society, 6(18):446–461, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Tizard:1954:NVS

[Tiz54] Henry Thomas Tizard. Nevil Vincent Sidgwick, 1873–1952. Obituary Notices of Fellows of the Royal Society, 9(1):236–258, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Toit:1948:AWR

[Toi48] Alexander Logie Du Toit. Arthur William Rogers, 1872–1946. Obituary Notices of Fellows of the Royal Society, 5(16):790–802, May 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Townsend:1943:JBP

[Tow43] John Sealy Edward Townsend. Jean Baptiste Perrin, 1870–1942. Obituary Notices of Fellows of the Royal Society, 4(12):301–305, November 1, 1943. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Turnbull:1938:EBE

[Tur38] Herbert Westren Turnbull. Edwin Bailey Elliott, 1851–1937. Obituary Notices of Fellows of the Royal Society, 2(6):424–431, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Turnbull:1941:AY

[Tur41] Herbert Westren Turnbull. Alfred Young, 1873–1940. Obituary Notices of Fellows of the Royal Society, 3(10):761–778, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Twyman:1938:WT

[Twy38] Frank Twyman. William Taylor, 1865–1937. Obituary Notices of Fellows of the Royal Society, 2(6):362–365, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Verney:1941:AJC

[VB41] Ernest Basil Verney and Joseph Barcroft. Alfred Joseph Clark, 1885– 1941. Obituary Notices of Fellows of the Royal Society, 3(10):969–984, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

W:1932:ETN

[W.32a] A. S. W. Edwin Tulley Newton, 1840–1930. Obituary Notices of Fellows of the Royal Society, 1(1):5–7, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

W:1932:WDM

[W.32b] D. M. S. W. William Diller Matthew, 1871–1930. Obituary Notices of Fellows of the Royal Society, 1(1):71–74, December 1, 1932. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

W:1934:FLK

[W.34] H. W. Finlay Lorimer Kitchin, 1870–1934. Obituary Notices of Fellows of the Royal Society, 1(3):349–352, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Watson:1940:GAB

[Wat40a] David Meredith Seares Watson. George Albert Boulenger, 1858–1937. Obituary Notices of Fellows of the Royal Society, 3(8):13–17, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Watts:1940:APC

[Wat40b] William Whitehead Watts. Arthur Philemon Coleman, 1852–1939. Obituary Notices of Fellows of the Royal Society, 3(8):117–122, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Watson:1950:CFC

[Wat50] D. M. S. Watson. Clive Forster-Cooper 1880–1947. Obituary Notices of Fellows of the Royal Society, 7(19):82–93, November 1, 1950. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Watson:1952:RB

[Wat52] David Meredith Seares Watson. Robert Broom 1866–1951. Obituary Notices of Fellows of the Royal Society, 8(21):36–70, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wenyon:1938:HSW

[Wen38] Charles Morley Wenyon. Henry Solomon Wellcome, 1853–1938. Obituary Notices of Fellows of the Royal Society, 2(6):229–238, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wenyon:1944:WY

[Wen44] Charles Morley Wenyon. Warrington Yorke 1883–1943. Obituary Notices of Fellows of the Royal Society, 4(13):523–545, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Weyl:1944:DH

[Wey44] Hermann Weyl. David Hilbert 1862–1943. Obituary Notices of Fellows of the Royal Society, 4(13):547–553, November 1, 1944. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Gray:1940:HMD

[WGS40] Robert Whytlaw-Gray and G. F. Smith. Harry Medforth Dawson, 1876–1939. Obituary Notices of Fellows of the Royal Society, 3(8):139– 154, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1935:HMM

[Whi35] E. T. Whittaker. Hector Munro Macdonald, 1865–1935. Obituary Notices of Fellows of the Royal Society, 1(4):550–558, December 1, 1935. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1936:ACD

[Whi36] E. T. Whittaker. Alfred Cardew Dixon, 1865–1936. Obituary Notices of Fellows of the Royal Society, 2(5):165–174, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whiddington:1939:VCW

[Whi39] Richard Whiddington. Victor Christian William Cavendish, the Duke of Devonshire, 1868–1938. Obituary Notices of Fellows of the Royal Society, 2(7):557–559, January 1, 1939. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1940:RAS

[Whi40] Edmund Taylor Whittaker. Ralph Allen Sampson, 1866–1939. Obituary Notices of Fellows of the Royal Society, 3(8):221–226, January 1, 1940. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1941:VV

[Whi41] Edmund Taylor Whittaker. Vito Volterra, 1860–1940. Obituary Notices of Fellows of the Royal Society, 3(10):691–729, December 1, 1941. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1942:ARF

[Whi42] E. T. Whittaker. Andrew Russell Forsyth, 1858–1942. Obituary Notices of Fellows of the Royal Society, 4(11):208–227, November 1, 1942. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1948:ANW

[Whi48] Edmund Taylor Whittaker. Alfred North Whitehead, 1861–1947. Obituary Notices of Fellows of the Royal Society, 6(17):280–296, November 1, 1948. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1949:PHC

[Whi49] Edmund T. Whittaker. Philip Herbert Cowell, 1870–1949. Obituary Notices of Fellows of the Royal Society, 6(18):375–384, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1951:AWC

[Whi51] Edmund Taylor Whittaker. Arthur William Conway, 1875–1950. Obituary Notices of Fellows of the Royal Society, 7(20):329–340, November 1, 1951. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whitehead:1952:EJC

[Whi52] John Henry Constantine Whitehead. Elie Joseph Cartan 1869–1951. Obituary Notices of Fellows of the Royal Society, 8(21):71–95, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Whittaker:1954:EWB

[Whi54] Edmund T. Whittaker. Ernest William Barnes, 1874–1953. Obituary Notices of Fellows of the Royal Society, 9(1):14–25, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wigglesworth:1949:ADI

[Wig49] Vincent Brian Wigglesworth. Augustus Daniel Imms, 1880–1949. Obituary Notices of Fellows of the Royal Society, 6(18):462–470, November 1, 1949. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wilson:1934:WCU

[Wil34] J. S. Wilson. William Cawthorne Unwin, 1838–1933. Obituary Notices of Fellows of the Royal Society, 1(3):167–178, December 1, 1934. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wilson:1938:SGE

[Wil38] James Thomas Wilson. Sir Grafton Elliot Smith, 1871–1937. Obituary Notices of Fellows of the Royal Society, 2(6):322–333, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wilson:1953:CHL

[Wil53] William Wilson. Charles Herbert Lees, 1864–1952. Obituary Notices of Fellows of the Royal Society, 8(22):522–528, November 1, 1953. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Wishart:1954:EPC

[Wis54] George M. Wishart. Edward Provan Cathcart, 1877–1954. Obituary Notices of Fellows of the Royal Society, 9(1):34–53, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Woodward:1936:HFO

[Woo36] Arthur Smith Woodward. Henry Fairfield Osborn, 1857–1935. Obituary Notices of Fellows of the Royal Society, 2(5):66–71, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Woodward:1938:WJS

[WW38] Arthur Smith Woodward and William Whitehead Watts. William Johnson Sollas, 1849–1938. Obituary Notices of Fellows of the Royal Society, 2(6):265–281, January 1, 1938. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Yates:1952:GUY

[Yat52] Frank Yates. George Udny Yule 1871–1951. Obituary Notices of Fellows of the Royal Society, 8(21):308–323, November 1, 1952. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Yule:1936:KP

[YF36] George Udny Yule and Louis Napoleon George Filon. Karl Pearson, 1857–1936. Obituary Notices of Fellows of the Royal Society, 2(5):72– 110, December 1, 1936. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).

Young:1954:JCD

[You54] Frank George Young. Jack Cecil Drummond 1891–1952. Obituary Notices of Fellows of the Royal Society, 9(1):98–129, November 1, 1954. CODEN ???? ISSN 1479-571X (print), 2053-9118 (electronic).